### v1

# Mixed-State Measurement-Induced Phase Transitions in Imaginary-Time Dynamics

Abstract
Mixed-state phase transitions have recently attracted growing attention as a new frontier in nonequilibrium quantum matter and quantum information. In this work, we introduce the measurement-dressed imaginary-time evolution (MDITE) as a novel framework to explore mixed-state quantum phases and decoherence-driven criticality. In this setup, alternating imaginary-time evolution and projective measurements generate a competition between coherence-restoring dynamics and decoherence-inducing events. While reminiscent of monitored unitary circuits, MDITE fundamentally differs in that the physics is encoded in decoherent mixed states rather than in quantum trajectories. We demonstrate that this interplay gives rise to a new class of mixed-state phase transitions, using numerical simulations of the one-dimensional transverse-field Ising model and the two-dimensional dimerized Heisenberg model. Furthermore, we provide a diagrammatic representation of the evolving state, which naturally enables efficient studies of MDITE with quantum Monte Carlo and other many-body numerical methods, thereby extending investigations of mixed-state phase transitions to large-scale and higher-dimensional Hamiltonians. Our results highlight MDITE as a powerful paradigm for investigating non-unitary dynamics and the fundamental role of decoherence in many-body quantum systems.
I Introduction
Decoherence plays a pivotal role in quantum physics, explaining the emergence of classical behavior from quantum systems through their inevitable interaction with the environment [Zeh1970decoherence, Zurek198pointer, Zurek1982superselection, Zurek2001EnvInduced, Zurek2003review, Schlosshauer2005review, cai2013algebraic, cai2014identifying, yan2018interacting, Schlosshauer2019review]. By entangling with environmental degrees of freedom, quantum superpositions decohere into statistical mixtures within a stable pointer basis, suppressing quantum interference and giving rise to effectively classical outcomes [Zurek198pointer, Zurek1982superselection]. Beyond its fundamental implications, decoherence also poses a major challenge in quantum information science, as it undermines superposition and entanglement—essential resources for achieving quantum advantage [NielsenChuang2010, Bennett1996purification, Bravyi2005universal, Gottesman2009errorcorrection, Terhal2015errorcorrection].
In contrast to uncontrollable environmental noise, projective quantum measurements provide a sharply defined and experimentally controllable way to introduce decoherence. This feature has made them a powerful tool for investigating open quantum systems. It has been recently discovered that, incorporating coherent dynamics into this setting leads to novel nonequilibrium phases and critical behaviors, collectively referred to as measurement-induced phase transitions (MIPTs) [liyaodong2018clifford, liyaodong2019clifford, SangSQ2021measure-protect, Skinner2019, Zabalo2022operator-scaling, fisher2023randomcircuit, Turkeshi2020measurement-criticality, Turkeshi2021from-infinite-to-zero-clicks, Lu2021spacetime-duality, Nahum2021measurement-entanglement, Sharma2022measurement-criticality, Szyniszewski2020universality-mpt, Tirrito2023fcs-mite, Zhou2023entanglement-criticality, Fux2024magiccircuit, Liushuo2024noise-mpt, Tsitsishvili2024mite-nonmarkovian, Chen2025subsystem, Chen2025mipt, zhu2023mipt]. The cannonical setting of MIPTs focuses on quantum trajectories [Wiseman1996quantum-trajectories, Plenio1998quantumjump, Daley2014quantumytrajectories]—ensembles of pure states conditioned on specific measurement outcomes—where the competition between scrambling unitary dynamics and disentangling measurements controls distinct scalings of entanglement.
A complementary, yet fundamentally distinct, perspective arises by considering the measurement-averaged mixed state, which directly captures the universal characterstics of decoherence [Garratt2023nonlocal-measure, Nahum2020majoranadecoherence, Lee2023decoh_criticality, Amin2024symmetry-mixed-ent, Passarelli2024mixed-mite, Su2024higherform-symmetry, Tantivasadakarn2024longrange-ent-measure, Fanruihua2024diagnostics, wangzhu2025decoherence]. In contrast to conventional MIPTs which are fundamentally rooted in quantum trajectories of monitored unitary circuits, we instead consider a mixed-state perspective based on the interplay between imaginary-time evolution (ITE) and projective measurements—two intrinsically non-unitary processes governing the dynamics. Building on this idea,, we introduce the measurement-dressed imaginary-time evolution (MDITE) (Sec. II).
Arising from a Wick rotation of the Schrödinger equation, ITE replaces real-time unitary operator $e^{-itH}$ with its non-unitary counterpart $e^{-\tau H}$, where $H$ is a Hamiltonian and $\tau$ denotes time. Unlike the oscillatory behavior of real-time quantum dynamics, ITE is intrinsically dissipative: high-energy excitations are exponentially suppressed so that generic initial states relax toward the ground state of the system described by $H$. Moreover, applying finite-time ITE to the maximally mixed state produces a thermal Gibbs state, highlighting its natural connection to finite-temperature statistical physics. While ITE was originally introduced as a mathematical tool for ground-state and thermal-state preparation, recent advances in quantum algorithms [Motta2020qite, McArdle2019vqite, Yuan2019theoryofvariational, Benedetti2021qite, Kondappan2023qite, Mao2023qite, Ding2024qite] and quantum hardware [Nishi2021qite-nisq, Cao2022qite, ZhangSX2024qite] have enabled its experimental realization. This also establishes the physical meaning of the MDITE we propose: on noisy quantum devices, the competition between ITE and decoherence arises naturally, providing a rich and realistic platform for studying novel many-body phenomena in open quantum systems.
Specifically, our MDITE setup begins with a maximally mixed state and alternates between ITE controlled by a many-body Hamiltonian and random projective measurements. In the absence of measurement, ITE acts as a deterministic mechanism, steering the system toward the ground state of the Hamiltonian. Over extended times, this evolution suppresses classical uncertainty while reinforcing quantum coherence. In contrast, nonzero projective measurements tend to amplify classical uncertainty and induce decoherence, counteracting the ordering effects of ITE. This competition suggests the existence of mixed-state phase transitions with respect to the stationary state of the process driven by the strength of decoherence, evolution time, and parameters in the Hamiltonian. As we will demonstrate in Sec. IV and V, through large-scale numerical simulations of the one-dimensional transverse-field Ising model and the two-dimensional columnar dimerized Heisenberg model, we reveal the existence of such mixed-state phase transitions, which can be characterized using conventional linear order parameters—such as spin magnetization—without relying on information-theoretic proxies.
It is important to emphasize that while our study focuses on a specific setup, the underlying framework is general and readily extends to other decoherence channels interacting with ITE. This establishes a broader paradigm of mixed-state phases and criticalities emerging from the interplay between imaginary-time dynamics and decoherence. Moreover, the rich diversity of ground states and exotic collective behaviors supported by different Hamiltonians significantly expands the scope for discovering novel many-body phenomena in open quantum systems. We will revisit the exciting potential for future exploration of this generalized paradigm in Sec. VI.
This paper is organized as follows. In Sec. II, we formally define the MDITE protocol and highlight its key features. In Sec. III, we provide a schematic representation of the MDITE process, which not only clarify the setup but also facilitate numerical studies including quantum Monte Carlo (QMC) methods. The one- and two-dimensional Hamiltonians studied in this work, along with the corresponding numerical results for mixed-state phases and criticalities, are introduced and analyzed in Secs. IV and V. Finally, in Sec. VI, we summarize the work and discuss several promising directions for future research.
II Protocol
In this section, we elaborate the measurement-dressed imaginary-time evolution (MDITE) protocol. While our discussion focuses on qubit systems, it can be readily generalized to systems with arbitrary local dimensions. As both the ITE and the projective measurements can be implemented on quantum circuits [Motta2020qite, McArdle2019vqite, Yuan2019theoryofvariational, Benedetti2021qite, Kondappan2023qite, Mao2023qite, Ding2024qite, Nishi2021qite-nisq, Cao2022qite, ZhangSX2024qite], we will describe the MDITE process in the language of quantum circuits.
Given a Hamiltonian $H$ describing a quantum many-body system of $N$ qubits, its associated MDITE protocol is defined as follows:
  1. 1.
    Input state: The protocol starts with the maximally mixed state $\rho_{0}=I_{2^{N}}/2^{N}$, where $I_{2^{N}}$ is the identity operator of dimension $2^{N}$.
  2. 2.
    Evolution: For each circuit layer $k=1,2,\dots$, the state evolves through two consecutive operations. First, a probabilistic projective measurement channel $\mathcal{M}_{p}$ is applied, where the measurement rate $p$ controls the probability that the qubits are measured. This is followed by an ITE with respect to $H$, i.e.,
    $\rho_{k}=\frac{1}{\mathcal{N}_{k}}e^{-\frac{\tau}{2}H}\mathcal{M}_{p}(\rho_{k-1})e^{-\frac{\tau}{2}H},$ (1)
    where $\mathcal{N}_{k}=\text{Tr}\!\left[e^{-\frac{\tau}{2}H}\mathcal{M}_{p}(\rho_{k-1})e^{-\frac{\tau}{2}H}\right]$ is the normalization factor ensuring $\text{Tr}(\rho_{k})=1$.
  3. 3.
    Output state: After applying $n_{d}$ layers, the final mixed state $\rho_{n_{d}}$ is obtained as the output.
In this study, we consider $\mathcal{M}_{p}$ to consist of probabilistic projective measurements in the computational basis. Specifically, each qubit is independently projectively measured with probability $p$, and left unchanged with probability (1p)(1-p). We are interested in the regime where, after a sufficiently large number of circuit layers $n_{d}$, the system approaches a stationary state, i.e., $\rho_{n_{d}}\approx\rho_{n_{d}+1}$. However, we stress that the existence of a stationary state is not guaranteed for general measurement channels. Nevertheless, for the examples considered in this work, such a stationary state indeed exists, as we will show in Secs. IV and V.
The dynamical process of MDITE involves a competition between two effects: the ITE operator $e^{-\frac{\tau}{2}H}$, which drives the system toward its ground state, and the measurement channel $\mathcal{M}_{p}$, which introduces decoherence and tends to produce a mixed state. As the computational basis is generally not the eigenbasis of $H$, ITE induces quantum fluctuations and reduces classical uncertainty. This suggests potential mixed-state phase transitions with respect to the stationary state $\rho_{n_{d}}$ when tuning the parameters in the protocol. It is important to note that the time step $\tau$ of ITE and the measurement rate $p$ serve as primary parameters controlling the behaviors of the stationary state $\rho_{n_{d}}$. Moreover, Hamiltonian parameters (e.g., coupling strength) also play a crucial role, resulting in a high-dimensional parameter space. This enriches the landscape of behaviors of the output state, rendering the setup of MDITE particularly compelling for investigating different novel many-body physics in mixed states.
III Diagrammatic representation of the monitored imaginary-time evolution
[…]
III.1 ITE without measurements
We begin by examining the case where no measurements are applied throughout the ITE processes, namely when the measurement rate is set to $p=0$. The state after the first circuit layer is
$\rho_{1}\propto e^{-\frac{\tau}{2}H}I_{2^{N}}e^{-\frac{\tau}{2}H}=e^{-\tau H},$ (2)
which can be written as
$|\rho_{1}=\sum_{s,s^{\prime}}\langle s^{\prime}|\rho_{1}|s\rangle\ket{s^{\prime}}\bra{s}$ (3)
[…]
Similarly, the output state after the second circuit layer, $\rho_{2}\propto e^{-2\tau H}$, is depicted in Fig. 1(b). We note that the intermediate states $|\ket{r^{\prime}}\bra{r}$ in the representation of $\rho_{2}$ [Fig. 1(b)] correspond to $|\ket{s^{\prime}}\bra{s}$ in the representation of $\rho_{1}$ [Fig. 1(a)], respectively. This follows from $\rho_{2}\propto e^{-\tau/2}\rho_{1}e^{-\tau/2}$, where the two ITE operators act on the upper and lower sides of $\rho_{1}$ in Fig. 1(a). In a manner analogous to tensor-network contraction, these operators propagate $\rho_{1}$, represented by $|\ket{r^{\prime}}\bra{r}$, into $\rho_{2}$, represented by $|\ket{s^{\prime}}\bra{s}$.
Notably, Fig. 1(a) and (b) can be interpreted as two ensembles, due to the summation over all possible $|\ket{s^{\prime}}\bra{s}$ in Eq. (3). We will revisit this notion of an ensemble in Sec. III.4. The cases of $\rho_{k}$ with $k>2$ follow in complete analogy, with their diagrammatic representations obtained recursively in the same way.
[…]
III.2 ITE with deterministic measurements
We next consider the case in which all qubits are measured deterministically in each circuit layer. Note that the output state $\rho_{1}\propto e^{-\tau H}$ after the first circuit layer is independent of the choice of measurement channel $\mathcal{M}_{p}$ and whether $\mathcal{M}_{p}$ is applied, since measurements on a maximally mixed state always leave it unchanged. Therefore, the diagrammatic representation of $\rho_{1}$ under deterministic measurements remains Fig. 1(a).
In the second circuit layer, we first obtain the measurement-averaged state after the measurement channel $\mathcal{M}_{p}$, which is
$|.\bar{\rho}_{1}=\sum_{s}\langle s|\rho_{1}|s\rangle\ket{s}\bra{s}$. (4)
[…]
After the ITE, the resulting state is $\rho_{2}\propto e^{-\tfrac{\tau}{2}H}\,\bar{\rho}_{1}\,e^{-\tfrac{\tau}{2}H}$, which is the output state after the second circuit layer. Analogous to the discussion in Sec. III.1, $\rho_{2}$ can be represented diagrammatically as in Fig. 2(b). Here, the state $|\ket{r}\bra{r}$ corresponds to the state $|\ket{s}\bra{s}$ in $\bar{\rho}_{1}$ shown in Fig. 2(a). It is straightforward to generalize the discussion to a general $\rho_{k}$ $(k>2)$, and the diagrammatic representations also follow recursively.
Figure 2: Diagrammatic representations of the MDITE under deterministic projective measurements. (a) Compared with Fig. 1, the measurement on $\rho_{1}$ identifies the bra and ket components of $\rho_{1}$ in Eq. (3), forming a closed loop for the measurement-averaged state $\bar{\rho}_{1}$. (b) The output state $\rho_{2}$ after the second circuit layer.
III.3 ITE with deterministic measurements on subsystem $A$ only
As the final preparatory step before discussing the general case of probabilistic measurements on each qubit, we consider the scenario in which all qubits in subsystem $A$ are measured projectively with probability one, while the complementary subsystem $B\equiv\overline{A}$ remains unmeasured.
[…]
where $s_{A}\in\{0,1\}^{\otimes N_{A}}$ and $s_{B}\in\{0,1\}^{\otimes N_{B}}$, with $N_{A}$ and $N_{B}$ denoting the number of qubits in subsystems $A$ and $B$, respectively.
[…]
Figure 3: Diagrammatic representation of the MDITE where only subsystem AA is projectively measured in a deterministic manner. The label “AA” (“BB”) indicates that the diagram applies to all qubits in subsystem AA ($B$). Since the computational basis is local, the evolutions of the reduced density matrices $(\rho_{0})_{A}=\text{Tr}_{B}(\rho_{0})$ and $(\rho_{0})_{B}=\text{Tr}_{A}(\rho_{0})$ can be depicted separately in (a) and (b), respectively.
III.4 ITE with probabilistic measurements
Now we extend the discussions above to a general case in which each site is independently measured with probability $p$ in each circuit layer.
For a system of $N$ qubits and $k$ circuit layers, the binary choice of measuring or not measuring each qubit in each layer yields $2^{(k-1)N}$ distinct ensembles, which together form an extended ensemble. In this sense, the examples in Figs. 1(b), 2(b), and 3 correspond to three of the $2^{N}$ possible ensembles when $k=2$.
The diagrammatic representation with probabilistic measurements now consists of $N$ components, a natural generalization of Fig. 3, one for each qubit. In Fig. 4, we present an additional example for $N=2$ and $k=3$, illustrating two of the $2^{4}$ possible ensembles.
[…]
III.5 QMC simulations
Importantly, the diagrammatic representation of the evolving state $\rho_{k}$ naturally facilitates QMC studies of $\rho_{n_{d}}$ based on imaginary-time path integrals (after applying $n_{d}$ layers, we obtain the final output state $\rho_{n_{d}}$), such as the stochastic series expansion (SSE) [Sandvik1992sse, Sandvik1999sse, Sandvik2010lecture, syljuaasen2002quantum, sandvik2019stochastic, Yan2019sweeping, yan2022global], as long as the Hamiltonian is sign-problem-free. In addition, it also makes it straightforward to implement tensor-network–based simulations [Ran2020tensor]. Therefore, we can study the stationary state $\rho_{n_{d}}$ by performing QMC simulations on the generalized partition function $\mathcal{Q}_{n_{d}}\propto\text{Tr}(\rho_{n_{d}})$, which is associated with the extended ensemble.
[…]
Therefore, we have
$\mathcal{Q}_{n_{d}}=\sum_{\{A_{k}\}}\prod_{l=2}^{n_{d}}\bigg[p^{N_{A_{l}}}(1-p)^{N-N_{A_{l}}}\bigg]Q_{n_{d}}(\{A_{k}\})$ (6)
where $N_{A_{l}}$ denotes the number of qubits in subsystem $A_{l}$. Since different ensembles $\{Q_{n_{d}}(\{A_{k}\})\}$ carry different weights determined by the measurement rate $p$, transitions between ensembles can be implemented in QMC simulations by designing appropriate selection probabilities that satisfy the detailed balance condition.
Within the SSE framework, we have developed an unbiased and efficient QMC algorithm to simulate Eq. (6), applicable to Hamiltonians for ITEs in arbitrary spatial dimensions. While the present work focuses on the stationary state $\rho_{n_{d}}$ ($n_{d}\to\infty$), the algorithm also enables studies of the dynamics of the MDITE process. Further details of the QMC algorithm are provided in Appendix A and B.
IV One-dimensional example with discrete symmetry
IV.1 Model and observables
As the first example, we consider the one-dimensional (1D) transverse-field Ising model (TFIM) [pfeuty1970one] as the Hamiltonian in the MDITE protocol, which is given by
$H=-\sum_{i}Z_{i}Z_{i+1}-h\sum_{i}X_{i}$ (7)
where $Z_{i}$ and $X_{i}$ are the Pauli operators acting on site $i$, and $h$ is the transverse field strength. The computational basis is chosen to be the $Z$-basis. The length of the chain is denoted by $L\equiv N$, and the periodic boundary condition, i.e., $N+1\equiv 1$, is considered.
This model has a global $\mathbb{Z}_{2}$ symmetry and the ground state exhibits a quantum phase transition at $h_{c,\mathrm{GS}}=1$, as determined by the Kramers-Wannier duality, which belongs to the 2D classical Ising universality class. For $h<h_{c,\mathrm{GS}}$, the system is in a ferromagnetic phase with nonzero magnetization, while for $h>h_{c,\mathrm{GS}}$, it is in a paramagnetic phase with zero magnetization.
Though free energy is not a well-defined observable in the MDITE protocol, the magnetization of spins remains accessible and physical, serving as an important proxy for characterizing the stationary state $\rho_{n_{d}}$. Specifically, for the TFIM, we examine the absolute magnetization $\langle|m|\rangle$, while also considering the second moment $\langle m^{2}\rangle$ and the fourth moment $\langle m^{4}\rangle$, from which the Binder ratio $R_{2}=\langle m^{4}\rangle/\langle m^{2}\rangle^{2}$ is constructed. Note that while the magnetization is defined as $\langle m\rangle:=\langle\sum_{i=1}^{N}Z_{i}\rangle$, $\langle|m|\rangle$ does not correspond to $\langle|\sum_{i=1}^{N}Z_{i}|\rangle$. Instead, in QMC simulations, $\langle|m|\rangle$ represents the ensemble average of the absolute magnetization measured in individual classical QMC samples [Sandvik2010lecture]. For the standard TFIM, observables $\langle|m|\rangle$ and $R_{2}$ are important for identifying the quantum critical point [blote2002cluster, huang2020worm, sandvik2003stochastic, zhou2022quantum, yan2023quantum]. In the ferromagnetic phase, the magnetization distribution is double-peaked around $\pm\langle|m|\rangle$, leading to $R_{2}\to 1$, while in the paramagnetic phase, it becomes Gaussian, yielding $R_{2}\to 3$. Remarkably, as we will show below, these observables can also be used to characterize the mixed-state phase transition of $\rho_{n_{d}}$ in the presence of measurements.
IV.2 Stationary state
First, we show the existence of the stationary state in the MDITE protocol, as noted in Sec. II. Fig. 5 shows that for the control parameters $(\tau,h,p)=(1,1.8,0.66)$—which we will later identify as a critical point—the state $\rho_{n_{d}}$ converges to a stationary state in the limit $n_{d}\to\infty$.
This is evidenced by the rapid convergence of both $\langle m^{2}\rangle$ and the Binder ratio $R_{2}$ with increasing circuit depth. We observe similar convergence behavior for other parameter choices. In practice, we find that a circuit depth of $n_{d}=2L/\tau$ is sufficient to reach the stationary state for the 1D TFIM.
[…]
Figure 5: When setting $(\tau,h,p)=(1,1.8,0.66)$ for the MDITE with the 1D TFIM: (a) Convergence of the second moment $\langle m^{2}\rangle$ with increasing circuit depth for various system sizes $L$. (b) Convergence of the Binder ratio $R_{2}$ with increasing circuit depth for various system sizes $L$.
IV.3 Measurement-induced phase transition
[…]
Figure 6: For $(\tau,h)=(1,1.8)$ in the MDITE of the 1D TFIM: (a) Binder ratios $R_{2}$ for different system sizes $L$ cross at $p_{c}\approx 0.667$; (b) The inset panel shows $\langle|m|\rangle$ as a function of $p$ for various $L$, while the main panel presents the finite-size scaling and data collapse of $\langle|m|\rangle$.
For $(h,p)=(2.5,0.5)$: (c) Binder ratios $R_{2}$ for different system sizes $L$ cross at $\tau_{c}\approx 0.265$; (d) $\langle|m|\rangle$ as a function of $\tau$ for various $L$ and data collapse of $\langle|m|\rangle$.
For $(\tau,p)=(1.2,0.8)$, (e) Binder ratios $R_{2}$ for different system sizes $L$ cross at $h_{c}\approx 1.84$; (f) $\langle|m|\rangle$ as a function of $h$ for various $L$ and data collapse of $\langle|m|\rangle$.
(g) and (h): Critical surfaces viewed from different angles. To facilitate the distinction of critical $p$ corresponding to different parameters on the critical surface, the figure employs a colorbar to map $p$ to different colors.
We explore the three-dimensional parameter space $(\tau,h,p)$ by first fixing $\tau$ and $h$, and then scanning the measurement rate $p$ to determine whether it induces a phase transition. Our extensive numerical calculations reveal that tuning $p$ indeed drives a transition for a wide range of $(\tau,h)$ values when $h>1$.
One example is shown for $(\tau,h)=(1,1.8)$ in Fig. 6(a), where a clear crossing of the Binder ratio $R_{2}$ across different system sizes $L$ identifies a critical point $p_{c}$. The asymptotic behavior of $R_{2}$—tending to 3 for $p<p_{c}$ and to 1 for $p>p_{c}$—provides strong evidence for a mixed-state phase transition analogous to that in the classical Ising model. This indicates a disordered mixed phase below $p_{c}$ and an ordered mixed phase above it.
At $p=0$, the stationary state corresponds to the ground state of the 1D TFIM, which is in the paramagnetic phase for $h=1.8>1$. The results in Fig. 6(a) thus demonstrate that sufficiently strong projective measurements can drive the system into a mixed ferromagnetic phase, polarized along the $z$-axis. In contrast to measurement-induced entanglement-entropy transitions that require quantum trajectory analysis [Skinner2019, buchhold2021effective], this transition is detectable using conventional linear observables without the need for information-theoretic diagnostics.
To identify the universality class of the critical point, we perform a finite-size scaling analysis [privman1990finite] on
$x_{L}=\frac{p-p_{c}}{p_{c}}L^{1/\nu},\quad y_{L}=\langle|m|\rangle L^{\beta/\nu}$ (8)
where $\nu$ and $\beta$ are the critical exponents related to the correlation length and the magnetization. From the data collapse, we extract $p_{c}\approx 0.667$, $\nu\approx 1.08$, $\beta\approx 0.43$, and $\beta/\nu\approx 0.40$, whose stability is reported in Table 1. The corresponding plots are presented in Fig. 6(b). We note that the critical exponents $\nu$ and $\beta$ are quite different from that of the equilibrium 2D or 3D classical Ising criticality, which means this mixed-state phase transition is not a standard Ising phase transition, but within a new universality class.
$L_{\min}$ $p_{c}$ $\nu$ $\beta$ $\beta/\nu$
112 0.6668(6) 1.073(5) 0.422(8) 0.393(5)
128 0.6667(7) 1.077(6) 0.424(9) 0.394(6)
144 0.6665(9) 1.079(7) 0.43(1) 0.395(8)
160 0.666(1) 1.084(8) 0.43(1) 0.397(96)
Table 1: For $(\tau,h)=(1,1.8)$, we perform a data collapse of $\langle|m|\rangle$ using system sizes $L=112,128,144,160,176,192$. To test the stability of the extracted exponents, we first discard the $L=112$ data, setting $L_{\min}=128$, and then progressively increase $L_{\min}$ to $160$ to retain only the three largest system sizes. The resulting stable values are $p_{c}\approx 0.667$, $\nu\approx 1.08$, $\beta\approx 0.43$, and $\beta/\nu\approx 0.40$.
IV.4 $\tau$- and $h$-induced phase transitions
In addition to the measurement-induced transition driven by tuning $p$, the stationary state exhibits mixed-state phase transitions when either the imaginary-time step $\tau$ or the field strength $h$ is varied, with the other parameters fixed. As an example, for $(h,p)=(2.5,0.5)$, the Binder ratio $R_{2}$ shows a clear crossing at $\tau_{c}\approx 0.265$, with a scaling analysis of $\langle|m|\rangle$ yielding $\nu\approx 1.19$ [Fig. 6(c, d), Table 2]. Similarly, for $(\tau,p)=(1.2,0.8)$, we identify a transition at $h_{c}\approx 1.84$, characterized by $\nu\approx 1.18$ and $\beta\approx 0.46$ [Fig. 6(e, f), Table 3].
Although the values of $\nu$ and $\beta$ differ from those of the $p$-driven transitions, the ratio $\beta/\nu\approx 0.40$ remains consistent within error bars across all cases examined. This suggests that $\beta/\nu$ could serve as a universal critical exponent for mixed-state phase transitions in the 1D TFIM under the MDITE protocol with $p\neq 0$. In the following section, we elucidate the mechanism underlying these transitions and demonstrate that they fall within a unique universality class.
$L_{\min}$ $\tau_{c}$ $\nu$ $\beta$ $\beta/\nu$
112 0.2651(1) 1.1873(2) 0.480(1) 0.404(1)
128 0.2652(1) 1.189(1) 0.480(2) 0.404(2)
144 0.2651(2) 1.190(2) 0.480(4) 0.403(3)
160 0.2652(4) 1.191(6) 0.480(7) 0.403(6)
Table 2: For $(h,p)=(2.5,0.5)$, we perform a data collapse of $\langle|m|\rangle$ using system sizes $L=112,128,144,160,176,192$. To test the stability of the extracted exponents, we first discard the $L=112$ data, setting $L_{\min}=128$, and then progressively increase $L_{\min}$ to $160$ to retain only the three largest system sizes. The resulting stable values are $\tau_{c}\approx 0.265$, $\nu\approx 1.19$, $\beta\approx 0.48$, and $\beta/\nu\approx 0.40$.
$L_{\min}$ $h_{c}$ $\nu$ $\beta$ $\beta/\nu$
112 1.8375(3) 1.181(5) 0.46(1) 0.39(1)
128 1.8373(7) 1.18(1) 0.46(1) 0.39(1)
144 1.837(1) 1.19(2) 0.46(1) 0.38(1)
160 1.835(2) 1.21(3) 0.46(1) 0.38(2)
Table 3: For $(\tau,p)=(1.2,0.8)$, we perform a data collapse of $\langle|m|\rangle$ using system sizes $L=112,128,144,160,176,192$. To test the stability of the extracted exponents, we first discard the $L=112$ data, setting $L_{\min}=128$, and then progressively increase $L_{\min}$ to $160$ to retain only the three largest system sizes. The resulting stable values are $h_{c}\approx 1.84$, $\nu\approx 1.18$, $\beta\approx 0.46$, and $\beta/\nu\approx 0.40$.
IV.5 Analysis of the criticalities
The key to understanding these mixed-state phase transitions lies in the dual role of the $Z$-basis: it is both the basis in which projective measurements are performed and the basis in which the ordered phase of the Hamiltonian is defined. Importantly, quantum coherence is generated solely by the off-diagonal terms, namely the transverse field, via the ITE operator $e^{-\frac{\tau}{2}H}$.
In the regime $h>1$ with short imaginary-time steps ($\tau h\ll 1$), frequent projective measurements at high $p$ interrupt the buildup of quantum coherence, producing dynamics reminiscent of the quantum Zeno effect [Misra1977zeno, Itano1990zeno, Koshino2005zeno] when the circuit is sufficiently deep. The transverse field, which drives quantum fluctuations in the $X$-basis, is repeatedly suppressed before correlations can form, as the state is frequently projected onto the $Z$-basis. With off-diagonal coherence effectively eliminated, the dynamics become dominated by the diagonal Ising terms, steering the system toward a ferromagnetic phase analogous to that of the classical Ising model. In contrast, when the transverse field is sufficiently strong or the evolution time is sufficiently long ($\tau h\gg 1$), quantum coherence emerges rapidly within the imaginary-time interval. Coherence is restored faster than it can be suppressed by measurements, preventing the development of ferromagnetic order even at $p=1$. Hence, no phase transition occurs in this regime by tuning the measurement rate $p$.
Between these two limits, the competition between quantum coherence and measurement-induced decoherence establishes a critical balance. Tuning the parameters, including $p$, $\tau$, and $h$, to this balance drives a mixed-state phase transition separating the ordered and disordered phases. In other words, the transitions driven by tuning $p$, $\tau$ or $h$ can be understood through the same mechanism. The distinction lies solely in which parameter modulates the competition between coherent evolution and measurement-induced decoherence.
Building upon the consistent critical exponent ratio $\beta/\nu\approx 0.40$ observed across all dynamical transitions in our simulations, and supported by the unified dynamical mechanism described above, we argue that the mixed-state phase transitions related to the 1D TFIM fall within a single universality class when $p>0$. This universality indicates the existence of a continuous critical surface in the extended parameter space $(\tau,h,p)$ separating the paramagnetic and ferromagnetic phases.
To visualize this structure, we construct a schematic phase diagram based on approximate transition points obtained from Binder ratio crossings in several system sizes, as shown in Fig. 6(g) and (h). Note that this surface serves as a qualitative guide to the phase boundary rather than a precise determination of the critical manifold. Importantly, Fig. 6(h) shows that the critical surface shifts toward larger $\tau$ as $h$ approaches $1$. This trend is expected: longer evolution times require weaker transverse fields to allow ferromagnetic order to emerge under strong measurements. Consequently, in the limit $h\to 1$, the critical surface is expected to extend infinitely along the $\tau$-axis, permitting a transition at any $\tau$ by tuning $p$. A similar reasoning applies in the opposite limit, $h\to\infty$ and $\tau\to 0$.
In addition, the emergence of long-range correlations leading to the Ising-like phase with increasing measurement rate can also be understood qualitatively from the perspective of QMC simulations: a higher measurement rate naturally generates larger clusters that connect a broader range of sites during the QMC updates. This provides a direct and intuitive picture of how measurements drive ordering in the system of TFIM, highlighting the use of the diagrammatic representations in Sec. III and the advantage of QMC in revealing the underlying mechanism. Further details are provided in Appendix. C.
V Two-dimensional example with continuous symmetry
V.1 Model and observables
As the second example, we consider the 2D $S=1/2$ antiferromagnetic columnar dimerized Heisenberg model (CDHM) on a $L\times L$ square lattice [Matsumoto2001qmcdhm]. The Hamiltonian is given by
$H=\sum_{\langle ij\rangle_{1}}\mathbf{S}_{i}\cdot\mathbf{S}_{j}+g\sum_{\langle ij\rangle_{g}}\mathbf{S}_{i}\cdot\mathbf{S}_{j}$ (9)
where $\mathbf{S}_{i}$ is the spin operator at site $i$, and $\langle ij\rangle_{1}$ and $\langle ij\rangle_{g}$ denote two different sets of nearest-neighbor pairs, as shown in Fig. 7. We also consider the local $Z$-basis as the computational basis, and the periodic boundary condition is chosen. The model has a global $SU(2)$ symmetry and the ground state exhibits a quantum phase transition at $g_{c,\mathrm{GS}}\approx 1.90951$ [Sandvik2010lecture, Matsumoto2001qmcdhm, Ma2018Anomalous], which belongs to the 3D $O(3)$ universality class. The critical point separates a Néel ordered phase ($g<g_{c,\mathrm{GS}}$) from a dimerized phase ($g>g_{c,\mathrm{GS}}$).
Figure 7: Square lattice of the 2D CDHM, where the red thick bonds denote couplings with strength $g$, while the remaining bonds have unit strength.
Due to the antiferromagnetic nature of the Heisenberg interactions, we consider the staggered magnetization $\langle m_{z}\rangle=\Big\langle\sum_{i}(-1)^{x_{i}+y_{i}}Z_{i}\Big\rangle$, where $x_{i}$ and $y_{i}$ denote the integer coordinates of site $i$ in the lattice. As in the TFIM, we compute the absolute magnetization $\langle|m_{z}|\rangle$, the second and fourth moments of the magnetization, and the Binder ratio $R_{2,z}$. Furthermore, for the ground state of the CDHM, the Binder ratio approaches $R_{2,z}\to 9/5$ in the Néel phase and $R_{2,z}\to 3$ in the dimerized phase [sandvik1997finite, wang2006high, ding2018engineering, Ma2018Anomalous].
V.2 Stationary state
As in the 1D TFIM case, the MDITE protocol in this case also drives the system toward a stationary state in the limit of large circuit depth. Fig. 8 illustrates this convergence for representative parameters $(\tau,g,p)=(1,3.5,0.5)$, where both the staggered magnetization $\langle|m_{z}|\rangle$ and the Binder ratio $R_{2,z}$ quickly approach their stationary values. In practice, we again find that a circuit depth of $n_{d}=2L/\tau$ is sufficient to reach the stationary state, and this choice is used in all subsequent QMC simulations for the CDHM.
[…]
Figure 8: When setting $(\tau,g,p)=(1,3.5,0.5)$ for the MDITE with the 2D CDHM: (a) Convergence of the squared magnetization $\langle m^{2}\rangle$ with increasing circuit depth for various system sizes $L$. (b) Convergence of the Binder ratio $R_{2,z}$ with increasing circuit depth for various system sizes $L$.
V.3 Mixed-state phase transitions and analysis
The parameter space of the stationary state in this model is characterized by $(\tau,g,p)$. Compared with the TFIM Hamiltonian, the CDHM Hamiltonian not only provides a two-dimensional example but also exhibits a global continuous $SU(2)$ symmetry. Interestingly, even at $g=0$, the ITE operator contains off-diagonal components $(S^{+}_{i}S^{-}_{j}+S^{-}_{i}S^{+}_{j})/2$. Increasing $g$ introduces both diagonal and off-diagonal terms, in sharp contrast to the TFIM case where tuning the Hamiltonian parameter $h$ only modifies the off-diagonal part.
Following the same strategy as in the TFIM case, we fix two of the parameters in $\{\tau,g,p\}$ and vary the third to search for mixed-state phase transitions. Taking the example of tuning $p$ at fixed $(\tau,g)=(1,3.5)$ in Fig. 9(a) and (b), we observe a clear crossing of the Binder ratio $R_{2,z}$ at $p_{c}\approx 0.354$, with $R_{2,z}\to 3$ for $p<p_{c}$ and $R_{2,z}\to 1$ for $p>p_{c}$. If $p=0$, the system is in the dimerized phase which has no order in the staggered magnetization. This indicates a mixed-state phase transition from a disordered phase to an AFM Ising-like ordered phase along the $z$-axis, driven by increasing the measurement strength. Although the Binder ratio exhibits Ising-like characteristic, similar to the case of TFIM, the extracted critical exponents in the case of 2D CDHM do not correspond to any known universality class including the 3D Ising universality class for a closed system, as shown in Table 4. Similar phenomena can also be observed when tuning $g$ and $\tau$, as shown in Fig. 9(c-f) and Tables 5 and 6.
Importantly, the critical exponents $\nu$ and $\beta$ also differ as we tune $p$, $\tau$, or $g$, but the ratio $\beta/\nu\approx 0.9$ across all cases remains consistent within error bars in our simulations, as shown in Tables 4, 5, and 6. Therefore, we also argue that these mixed-state phase transitions in the 2D CDHM belong to a single universality class when $p>0$, despite the differences in $\nu$ and $\beta$.
To interpret these results, note that any nonzero measurement rate $p$ explicitly breaks the $SU(2)$ symmetry, favoring the $z$-axis. Although increasing $g$ in the Hamiltonian introduces both diagonal and off-diagonal terms, only the diagonal components are compatible with the projective measurements along the $z$-axis, which correspond to the AFM Ising interactions. When $g>g_{c,\mathrm{GS}}$, and the imaginary-time steps are short ($\tau g\ll 1$), a high measurement rate $p$ suppresses the off-diagonal coherence generated by the ITE operator, leading to dynamics dominated by the diagonal AFM Ising terms. This drives the system into an ordered phase with nonzero staggered magnetization along the $z$-axis. In contrast, if $\tau g\gg 1$, the off-diagonal coherence builds up rapidly, and the final stationary state remains disordered even at $p=1$. Therefore, for projective measurements along the $z$-axis, the MDITE with the 2D CDHM effectively generalizes the physics of the 1D TFIM to two dimensions.
Since the CDHM can also be simulated using cluster updates within QMC, the emergence of long-range correlations induced by measurements can naturally be understood qualitatively through the cluster formation mechanism, as discussed in Appendix. C. We emphasize that it would be interesting to explore other types of decoherence channels that preserve the $U(1)$ or $SU(2)$ symmetry, which we leave for future work. In such cases, more advanced update algorithms beyond conventional cluster methods for QMC stimulations would be necessary, potentially revealing qualitatively different types of phase transitions. Physically, the enhancement of long-range order induced by measurement can be interpreted as a suppression of quantum fluctuations—i.e., the off-diagonal elements of the density matrix—caused by the diagonal measurement employed in our setup.
The schematic critical surface associated with this model is shown in Fig. 9(g) and (h), which provides a qualitative view to understand the phase transition driven by $p$, $g$ and $\tau$. This critical surface also shows that large $g$ requires small evolution time, while long evolution time requires weak coupling $g$. This inverse relationship between $g$ and $\tau$ closely resembles that of the 1D TFIM.
[…]
Figure 9: For $(\tau,g)=(1,3.5)$ in the MITE of the 2D CDHM: (a) Binder ratios $R_{2,z}$ for different system sizes $L$ cross at $p_{c}\approx 0.354$. (b) The inset panel shows $\langle|m_{z}|\rangle$ as a function of $p$ for various $L$, while the main panel presents the finite-size scaling and data collapse of $\langle|m_{z}|\rangle$.
For $(g,p)=(3,0.1)$: (c) Binder ratios $R_{2,z}$ for different system sizes $L$ cross at $\tau_{c}\approx 0.468$; (d) $\langle|m|\rangle$ as a function of $\tau$ for various $L$ and data collapse of $\langle|m|\rangle$.
For $(\tau,p)=(2,0.3)$: (e) Binder ratios $R_{2,z}$ for different system sizes $L$ cross at $g_{c}\approx 2.8$; (f) $\langle|m|\rangle$ as a function of $g$ for various $L$ and data collapse of $\langle|m|\rangle$.
(g) and (h): Critical surfaces viewed from different angles. To facilitate the distinction of critical $p$ corresponding to different parameters on the critical surface, the figure employs a colorbar to map $p$ to different colors.
$L_{\min}$ $p_{c}$ $\nu$ $\beta$ $\beta/\nu$
$24$ 0.3537(4) 0.714(7) 0.650(11) 0.910(7)
$28$ 0.3537(4) 0.716(6) 0.651(11) 0.909(8)
$32$ 0.3538(5) 0.717(6) 0.650(12) 0.906(9)
Table 4: We set $(\tau,g)=(3.5,1)$ to extract the critical exponent of $\langle|m|\rangle$ in the MIPT by using system size $L=24,28,32,40,48$. To test the stability of the extracted exponent, we gradually discard the small system size to perform the data collapse. $L_{\min}$ is the smallest system size used in the data collapse from 24 to 32. The resulting stable values are $p_{c}\approx 0.354$, $\nu\approx 0.714$, $\beta\approx 0.65$, and $\beta/\nu\approx 0.90$.
$L_{\min}$ $\tau_{c}$ $\nu$ $\beta$ $\beta/\nu$
$24$ 0.4679(5) 0.799(6) 0.718(9) 0.898(5)
$28$ 0.4679(6) 0.801(6) 0.721(10) 0.900(7)
$32$ 0.4678(7) 0.801(6) 0.720(12) 0.899(9)
Table 5: We set $(g,p)=(3,0.1)$ to extract the critical exponent of $\langle|m|\rangle$ in the MIPT by using system size $L=24,28,32,40,48$. To test the stability of the extracted exponent, we gradually discard the small system size to perform the data collapse. $L_{\min}$ is the smallest system size used in the data collapse from 24 to 32. The resulting stable values are $\tau_{c}\approx 0.4679$, $\nu\approx 0.8$, $\beta\approx 0.72$, and $\beta/\nu\approx 0.90$.
$L_{\min}$ $g_{c}$ $\nu$ $\beta$ $\beta/\nu$
$24$ 2.806(3) 0.736(18) 0.654(26) 0.89(2)
$28$ 2.806(2) 0.739(17) 0.653(26) 0.88(2)
$32$ 2.806(2) 0.741(15) 0.650(26) 0.88(2)
Table 6: We set $(\tau,p)=(2,0.3)$ to extract the critical exponent of $\langle|m|\rangle$ in the MIPT by using system size $L=24,28,32,40,48$. To test the stability of the extracted exponent, we gradually discard the small system size to perform the data collapse. $L_{\min}$ is the smallest system size used in the data collapse from 24 to 32. The resulting stable values are $g_{c}\approx 2.806$, $\nu\approx 0.74$, $\beta\approx 0.65$, and $\beta/\nu\approx 0.90$.
VI Conclusion and discussions
In this work, we have introduced MDITE as a versatile framework for studying novel mixed-state phases and phase transitions in open quantum systems. By combining ITE with probabilistic projective measurements, MDITE enables the exploration of fully non-unitary dynamics and their stationary states, providing both conceptual insight and numerical accessibility for studying high-dimensional open quantum systems.
As illustrative examples, we applied MDITE to the 1D TFIM and the 2D CDHM, uncovering distinct mixed-state phases and associated phase transitions in the stationary states of the protocol. Remarkably, the Ising-like universalities observed in both 1D and 2D models are characterized by the ratio of critical exponents $\beta/\nu$, rather than the individual exponents. To the best of our knowledge, these represent new universality classes in open quantum systems. We provided a visualized understanding of these criticalities and further showed that the transitions can be interpreted in terms of cluster formation within the QMC simulations. These results demonstrate that the MDITE is an excellent framework for exploring nontrivial mixed states, yielding fresh insights into the physics of open quantum systems.
At the methodological level, we introduced a diagrammatic representation of the MDITE process, which facilitates QMC and other numerical simulations on a broad class of many-body Hamiltonians, including those exhibiting spontaneous symmetry breaking, topological phase transitions, and topologically ordered phases. While our analysis focused on magnetization and the Binder ratio, the computation of information-theoretic quantities—such as entanglement entropy [hasting2010renyi, stephan2012renyi, melko2010renyi, jonathan2020renyi, Zhao2022renyi, ding2024reweight, wangz2025reweight, tarabunga2025bell, yan2023unlocking, li2024relevant, mao2025sampling, wang2025extracting] and Rényi negativity (including entanglement witnesses based on partitial transpose moments) [ding2025tracking, Wukh2020negativity, Alba2013negativity, Wangfh2025negativity, chungchiamin2014negativity, yu2021pt, elben2020pt, Neven2021pt, tarabunga2025pt]—is straightforward in combination with recent QMC techniques.
Our work opens several promising avenues for future research. One direction is to further characterize the universality classes associated with the criticalities observed in our simulations. Moreover, our QMC approach allow systematic investigations of dynamical properties and nonequilibrium phenomena of the MDITE process. Another is to extend the framework to more complex Hamiltonians and general decoherence channels beyond projective measurements, thereby accessing a broader landscape of many-body phenomena. In particular, long-range entanglement or topological order in mixed states [lu2023charac, lu2023longrange, Fanruihua2024diagnostics, Lu2024disentangling, lessa2025higherformanomaly], as well as the strong-to-weak spontaneous symmetry-breaking [Lee2023decoh_criticality, Lessa2025swssb, sala2024swssb, gu2024swssb, huang2025swssb, zack2025renyi1corr, guo2025swssb, guoyc2025swssb, guoyc2025ldpo, zhang2025swssb] can be explored within the MDITE framework.
On the experimental side, MDITE could be implemented in platforms such as Rydberg atom arrays [Wu2021rydbergreview, Evered2023rydberg], superconducting qubits [Wendin2017superconduct], or trapped-ion systems [haffner2008trapped]. Such realizations would enable direct observation of the mixed-state criticalities predicted here, bridging theory and experiment.
Acknowledgements
We thank G.-Y. Zhu, B.-B. Miao and Z. Bi for helpful discussions. Zenan Liu thanks the China Postdoctoral Science Foundation under Grants No.2024M762935 and NSFC Special Fund for Theoretical Physics under Grants No.12447119. Zhe Wang thanks the China Postdoctoral Science Foundation under Grants No.2024M752898. This project is supported by the Scientific Research Project (No.WU2024B027) and the Start-up Funding of Westlake University. The authors thank the high-performance computing center of Westlake University and the Beijing PARATERA Tech Co.,Ltd. for providing HPC resources.
Appendix A A brief review of the stochastic series expansion quantum Monte Carlo
The stochastic series expansion (SSE) method is a versatile and efficient quantum Monte Carlo approach for investigating both finite-temperature and ground-state properties of spin and bosonic systems. It reformulates the partition function $Z=\text{Tr}(e^{-\beta H})$ by performing a Taylor expansion of the exponential operator. The expansion is carried out in a chosen basis $\{|\alpha\rangle\}$, enabling efficient stochastic sampling of physical observables.
By writing the Hamiltonian as a sum of local operators,
$H=-\sum_{a,b}H_{a,b},$ (10)
where $a$ refers to the type of operators (diagonal or off-diagonal operator; site or bond operator) and $b$ marks the spatial indices (e.g., the labels of sites or bonds), we have
$Z=$ $\sum_{\alpha}\langle\alpha|e^{-\beta H}|\alpha\rangle$
$=$ $\sum_{\alpha}\sum_{\{H_{a,b}\}}\frac{\beta^{n}(M-n)!}{M!}\langle\alpha|\prod^{M}_{i=1}H_{a(i),b(i)}|\alpha\rangle,$ (11)
where $M$ is the truncation of the expansion order. To ensure that all operator sequences $\prod_{i=1}^{M}H_{a(i),b(i)}$ have the same fixed length $M$, we have introduced a null operator $H_{-1,-1}$ serves as an identity element filling the unoccupied positions in each sequence.
In the SSE method, each Monte Carlo step consists of a diagonal update and an off-diagonal update. The diagonal update alternates between diagonal operators and null operators while preserving the detailed balance condition. Through this process, the total number of non-null operators $n$ in the operator sequence can change.
  1. (a)
    For each null operator, we try inserting a new diagonal operator at a random position with probability
    $P_{\mathrm{add}}(n\rightarrow n+1)=\min\bigg\{\frac{\beta N_{b}\langle\alpha|H_{b}|\alpha\rangle}{M-n},1\bigg\},$ (12)
    where $N_{b}$ denotes the total number of spatial bonds (or sites) available for insertion.
  2. (b)
    For each diagonal operator, we remove it with probability
    $P_{\mathrm{remove}}(n+1\rightarrow n)=\min\bigg\{\frac{M-n+1}{\beta N_{b}\langle\alpha|H_{b}|\alpha\rangle},1\bigg\}$ (13)
The off-diagonal update will change the number of operators and update between diagonal operators and off-diagonal operators. To improve the update efficiency, some nonlocal update scheme such as the loop update, cluster update, or directed loop update are typically used [Sandvik2010lecture, sandvik2003stochastic, Anders2002Quantum].
Next, we take the TFIM in Sec. IV as an example to illustrate the SSE method, and it is similar to discuss the CDHM in Sec. V. The Hamiltonian of the TFIM is given by
$H=-J\sum_{\langle ij\rangle}Z_{i}Z_{j}-h\sum_{i}X_{i}$ (14)
To express this Hamiltonian in the form of Eq. (10), we introduce the following operators:
$H_{0,i}$ $\equiv hI_{i}$
$H_{1,i}$ $\equiv hX_{i}$
$H_{2,\langle ij\rangle}$ $\equiv J(Z_{i}Z_{j}+I_{i}I_{j})$ (15)
In the computational basis, defined as the eigenbasis of the Pauli-$Z$ operators, $H_{0,i}$ represents a diagonal on-site operator associated with site $i$, $H_{1,i}$ is an off-diagonal on-site operator, and $H_{2,\langle ij\rangle}$ is a diagonal bond operator associated with the site pair $\langle ij\rangle$. The nonzero matrix elements of the operators defined in Eq. (A) are
$\langle\uparrow_{i}|H_{0,i}|\uparrow_{i}\rangle$ $=\langle\downarrow_{i}|H_{0,i}|\downarrow_{i}\rangle=h$
$\langle\uparrow_{i}|H_{1,i}|\downarrow_{i}\rangle$ $=\langle\downarrow_{i}|H_{1,i}|\uparrow_{i}\rangle=h$
$\langle\uparrow_{i}\uparrow_{j}|H_{2,\langle ij\rangle}|\uparrow_{i}\uparrow_{j}\rangle$ $=\langle\downarrow_{i}\downarrow_{j}|H_{2,\langle ij\rangle}|\downarrow_{i}\downarrow_{j}\rangle=2J$ (16)
[…]
Figure 10: The diagram for cluster update in the TFIM model. White rectangle and blue rectangle denotes the diagonal operator and site operator respectively. The yellow area represents a constructed cluster, with the purple solid line indicating the update line. The line stops upon encountering a site operator; when it meets a bond operator, it continues to extend branches from the other legs of that operator. The dashed lines represent periodic boundary conditions in imaginary time.
For the cluster update, we construct clusters according to the following two rules: (i) A cluster is terminated by a site operator; (ii) The update line continues to extend from the other legs of a bond operator, while each bond operator belongs to exactly one cluster. The process is repeated until all clusters are identified. An illustration of the cluster construction is shown in Fig. 10, where the yellow area is a cluster. Finally, each cluster is flipped with probability $1/2$. As we will show below, this update can be readily adapted to the simulation of the extended ensemble in Eq. (6).
Appendix B Quantum Monte Carlo algorithm for simulating the generalized partition function $\mathcal{Q}_{n_{d}}$
We still take the TFIM as an example. Compared with the standard partition function $Z$, the simulation of $\mathcal{Q}_{n_{d}}$ involves multiple replicas and probabilistic boundary conditions that connect them. Fig. 11 shows the SSE configuration of $\bar{\rho}_{1}\propto e^{-\tau H}$, whose QMC simulation is equivalent to that of a Gibbs state with partition function $Z=\text{Tr}(e^{-\tau H})$.
[…]
Figure 11: The measurement-averaged state $\overline{\rho}_{1}$ corresponds to the sampling configuration in the SSE stimulation. In the diagram, white rectangles denote diagonal operators, blue rectangles represent site operators, and red rectangles indicate auxiliary identity operators. Additional auxiliary identity operators (red rectangles) are introduced at each local site. Those auxiliary operators do not affect the physical properties of the system, and they facilitate the treatment of boundary conditions in simulations practically.
[…]
Figure 12: The measurement-averaged state $\overline{\rho}_{1}$ corresponds to the sampling configuration in the SSE stimulation, taking two circuit layers as an example. White rectangle and blue rectangle represents the diagonal operator and site operator respectively, while red rectangle denotes the auxiliary identity operators. In (a), the simulation performs a measurement operation after a time interval of $\tau/2$. At this moment, the manifold closes, and the two circuit layers are adhered together. In (b), no measurement is performed at $\tau/2$, which the manifold remains open, and the two circuits form a single larger layer. The dashed lines represent the periodic boundary conditions in imaginary time. The two subfigures correspond respectively to the two extreme cases of measurement rates $p=1$ and $p=0$.
Similarly, when the measurement rate is $p=1$, as discussed in Sec. III.2, the SSE configuration related to $\text{Tr}(\rho_{2})$ is illustrated in Fig. 12(a). Since the state undergoes two evolutions, the diagram contains two replicas of $e^{-\tau H}$. To perform the QMC simulations, we note that the boundary conditions must be consistent with the diagrammatic representation introduced in Sec. III, which imposes constraints on the spin configurations near the boundaries of each replica. Since the diagonal update in SSE is local and does not alter the spin states, it can be implemented in the same manner as that in a standard partiton function $Z$. For the nonlocal update, the boundary conditions require that the spins at the interfaces between different replicas remain consistent in Fig. 12(a). Consequently, when a cluster extends across an interface, it may continue into another replica. In practice, this is implemented using auxiliary identity operators: when two replicas are connected, their corresponding auxiliary identity operators are required to belong to the same cluster.
When the measurement rate is zero, the ensemble reduces to the standard partition function with a doubled inverse temperature, as illustrated in Fig. 12(b). In this case, the auxiliary operators are not involved in the cluster update.
As discussed in Sec. III.4, for a general measurement rate $p$, we only need to switch the simulated ensemble between Fig. 11(a) and (b) when the circuit layer is two. Specifically, before each Monte Carlo step, we consider the following merge-split process:
  1. (1)
    If two spins near the boundary in the corresponding replicas are in the same state but are not connected in the diagrammatic representation, they are merged with probability
    $P_{\mathrm{merge}}=\min\bigg\{\frac{p}{1-p},1\bigg\}$ (17)
  2. (2)
    If two spins are in the same state and already connected in the diagrammatic representation, they are split with probability
    $P_{\mathrm{split}}=\min\bigg\{\frac{1-p}{p},1\bigg\}$ (18)
The merge-split process can be naturally generalized to the case when the circuit layer is greater than two, i.e., more replicas of $e^{-\tau H}$, following the dicussions in Sec. III.5. For the Heisenberg model, the diagonal update, merge–split process, and nonlocal update are implemented in a manner analogous to those used for the TFIM under measurement. The cluster update scheme can be employed to perform updates between diagonal and off-diagonal operators in the Heisenberg model.
Appendix C Cluster formations and long-range correlations
In this section, from the perspective of QMC simulations, we provide an intuitive understanding of how measurements induce long-range correlations in the system, leading to the emergence of a mixed and ordered phase.
As introduced in Appendix. A, the nonlocal cluster update is a crucial step to ensure the ergodicity of the QMC algorithm for the TFIM. Therefore, larger clusters automatically formed during the update process correspond to stronger correlations between different sites. For example, in the 1D TFIM, when the transverse field is very strong $(h\to\infty)$, spins are dominated by the external field, and cluster updates produce only small, local clusters. Conversely, when the field is very weak, $(h\to 0)$, the Ising interactions dominate, connecting all sites into a single large clusters. This extensive connectivity generates strong correlations throughout the system, corresponding to the ferromagnetic phase with long-range order.
According to the merge-split process introduced in Appendix. B, the presence of measurements allows update lines to connect across replicas, effectively merging clusters from different replicas. Therefore, even starting with the paramagnetic ground state of the TFIM, measurements can induce the formation of larger clusters during the QMC updates, which is similar to increasing the Ising coupling strength for the ground state of the TFIM. This is the reason why measurements can drive the system into a mixed-ordered phase with long-range correlations, from the perspective of cluster formations in QMC simulations. Similar reasoning also applies to the CDHM, where a cluster update is also used.
For general Hamiltonians, the specific form of the nonlocal update scheme may differ from the cluster update employed in the TFIM and CDHM. Depending on the model, the corresponding update mechanism could also reveal new insights into the intrinsic properties of the states in the MDITE process.