### v3

# Mixed-State Measurement-Induced Phase Transitions in Imaginary-Time Dynamics

Abstract
Mixed-state phase transitions have recently attracted growing attention as a new frontier in nonequilibrium quantum matter and quantum information. In this work, we introduce the measurement-dressed imaginary-time evolution (MDITE) as a novel framework to explore mixed-state quantum phases and decoherence-driven criticality. In this setup, alternating imaginary-time evolution and projective measurements generate a competition between coherence-restoring dynamics and decoherence-inducing events. While reminiscent of monitored unitary circuits, MDITE fundamentally differs in that the physics is encoded in decoherent mixed states rather than in quantum trajectories. Using numerical simulations of the one-dimensional transverse-field Ising model and the two-dimensional columnar dimerized Heisenberg model, we demonstrate the existence of this kind of mixed-state phase transitions. Notably, these transitions appear to exhibit critical behavior inconsistent with known universality classes. In addition, we provide a diagrammatic representation of the evolving state, which naturally enables efficient studies of MDITE with quantum Monte Carlo and other many-body numerical methods, thereby extending investigations of mixed-state phase transitions to large-scale and higher-dimensional systems. Our results establish MDITE as a versatile platform for investigating mixed-state criticality and uncover new classes of decoherence-driven nonequilibrium phase transitions.
I Introduction
Decoherence plays a pivotal role in quantum physics, explaining the emergence of classical behavior from quantum systems through their inevitable interaction with the environment [123, 131, 132, 71, 133, 86, 7, 8, 115, 87]. By entangling with environmental degrees of freedom, quantum superpositions decohere into statistical mixtures within a stable pointer basis, suppressing quantum interference and giving rise to effectively classical outcomes [131, 132]. Beyond its fundamental implications, decoherence also poses a major challenge in quantum information science, as it undermines superposition and entanglement—essential resources for achieving quantum advantage [68, 3, 5, 28, 98].
In contrast to uncontrollable environmental noise, projective quantum measurements provide a sharply defined and experimentally controllable way to introduce decoherence. This feature has made them a powerful tool for investigating open quantum systems. It has been recently discovered that, incorporating coherent dynamics into this setting leads to novel nonequilibrium phases and critical behaviors, collectively referred to as measurement-induced phase transitions (MIPTs) [47, 48, 84, 89, 122, 25, 103, 102, 51, 65, 88, 94, 100, 119, 26, 49, 101, 11, 10, 129]. The cannonical setting of MIPTs focuses on quantum trajectories [111, 73, 16]—ensembles of pure states conditioned on specific measurement outcomes—where the competition between scrambling unitary dynamics and disentangling measurements controls distinct scalings of entanglement.
A complementary, yet fundamentally distinct, perspective arises by considering the measurement-averaged mixed state, which directly captures the universal characterstics of decoherence [27, 66, 43, 63, 70, 90, 95, 24, 106, 30, 31, 32, 130]. In contrast to conventional MIPTs which are fundamentally rooted in quantum trajectories of monitored unitary circuits, we instead consider a mixed-state perspective based on the interplay between imaginary-time evolution (ITE) and projective measurements—two intrinsically non-unitary processes governing the dynamics. Building on this idea,, we introduce the notion of measurement-dressed imaginary-time evolution (MDITE) (Sec. II).
Arising from a Wick rotation of the Schrödinger equation, ITE replaces real-time unitary operator $e^{-itH}$ with its non-unitary counterpart $e^{-\tau H}$, where $H$ is a Hamiltonian and $\tau$ denotes time. Unlike the oscillatory behavior of real-time quantum dynamics, ITE is intrinsically dissipative: high-energy excitations are exponentially suppressed so that generic initial states relax toward the ground state of the system described by $H$. Moreover, applying finite-time ITE to the maximally mixed state produces a thermal Gibbs state, highlighting its natural connection to finite-temperature statistical physics. While ITE was originally introduced as a mathematical tool for ground-state and thermal-state preparation, recent advances in quantum algorithms [64, 60, 121, 2, 41, 58, 20] and quantum hardwares [69, 9, 125] have enabled its experimental realization. This also establishes the physical meaning of the MDITE we propose: on noisy quantum devices, the competition between ITE and decoherence arises naturally, providing a rich and realistic platform for studying novel many-body phenomena in open quantum systems.
Our MDITE setup begins with a maximally mixed state and alternates between ITE controlled by a local Hamiltonian and random projective measurements. In the absence of measurement, ITE acts as a deterministic mechanism, steering the system toward the ground state of the Hamiltonian. Over extended times, this evolution suppresses classical uncertainty while reinforcing quantum coherence. In contrast, nonzero projective measurements tend to amplify classical uncertainty and induce decoherence, counteracting the ordering effects of ITE. This competition suggests the existence of mixed-state phase transitions with respect to the steady state of the process driven by the strength of decoherence, evolution time, and parameters in the Hamiltonian.
As we will demonstrate in Secs. IV, V and VI, through large-scale numerical simulations of the one-dimensional transverse-field Ising model and the two-dimensional columnar dimerized Heisenberg model, we reveal the existence of such mixed-state MIPTs. Remarkably, these transitions can be characterized by conventional linear order parameters, such as spin magnetization, without relying on information-theoretic diagnostics typically used in MIPTs based on quantum trajectories [47, 48, 84, 89, 122, 25, 103, 102, 51, 65, 88, 94, 100, 119, 26, 49, 101, 11, 10, 129]. Supported by theoretical discussions and numerical analysis of the critical behavior, our results indicate that, to the best of our knowledge, these transitions do not belong to any previously established universality class.
It is important to emphasize that while our study focuses on a specific setup, the underlying framework is general and readily extends to other decoherence channels interacting with ITE. This establishes a broader paradigm of mixed-state phases and criticalities emerging from the interplay between imaginary-time dynamics and decoherence. Moreover, the rich diversity of ground states and exotic collective behaviors supported by different Hamiltonians significantly expands the scope for discovering novel many-body phenomena in open quantum systems. We will revisit the exciting potential for future exploration of this generalized paradigm in Sec. VII.
This paper is organized as follows. In Sec. II, we introduce the MDITE protocol and discuss its key features. In Sec. III, we present a diagrammatic representation of the MDITE process, which clarifies the setup and enables efficient numerical approaches, including quantum Monte Carlo (QMC). Within this framework, the mixed-state MIPTs admit a natural interpretation in terms of cluster formation in the QMC updates. The one- and two-dimensional Hamiltonians studied in this work, together with the corresponding results for mixed-state phases and critical behavior, are presented in Secs. IV and VI. In particular, Sec. V provides important theoretical insights by examining some continuous limits in which the dynamics map onto effective equilibrium problems in a doubled Hilbert space, thereby elucidating the underlying mechanism of the transitions we discovered qualitatively. Finally, in Sec. VII, we summarize our findings and outline several directions for future research.
II Protocol
In this section, we formally define the measurement-dressed imaginary-time evolution (MDITE) protocol. While our discussion focuses on qubit systems, it can be readily generalized to systems with arbitrary local dimensions.
Given a Hamiltonian $H$ describing a quantum many-body system of $N$ qubits, its associated MDITE protocol is defined as follows:
  1. 1.
    Input state: The protocol starts with the maximally mixed state $\rho_{0}=I_{2^{N}}/2^{N}$, where $I_{2^{N}}$ is the identity operator of dimension $2^{N}$.
  2. 2.
    Evolution: For each discrete time step $k=1,2,\dots$, the state evolves through two consecutive operations:
    • A probabilistic projective measurement channel $\mathcal{E}_{p}$ in the computational basis is applied, where the measurement rate $p$ controls the probability that the qubits are measured;
    • An ITE with respect to $H$.
    The resulting state is
    $\rho_{k}=\frac{1}{\mathcal{N}_{k}}e^{-\frac{\tau}{2}H}\mathcal{E}_{p}[\rho_{k-1}]e^{-\frac{\tau}{2}H},$ (1)
    where $\mathcal{N}_{k}=\text{Tr}\!\left(e^{-\frac{\tau}{2}H}\mathcal{E}_{p}[\rho_{k-1}]e^{-\frac{\tau}{2}H}\right)$ is the normalization factor ensuring $\text{Tr}(\rho_{k})=1$.
  3. 3.
    Output state: After $n_{d}$ steps, the final mixed state $\rho_{n_{d}}$ is obtained as the output.
Specifically, we consider that each qubit is independently measured with probability $p$, and left unchanged with probability (1p)(1-p). Mathematically,
$\mathcal{E}_{p}=\prod_{i=1}^{N}\mathcal{E}_{p,i},$ (2)
with
$\mathcal{E}_{p,i}[\rho]\equiv$ $(1-p)\rho$ (3)
[…]
Here $|\ket{0}\equiv\ket{\uparrow}$ and $|\ket{1}\equiv\ket{\downarrow}$ denote the eigenstates of the local Pauli operator $Z_{i}$ acting on the $i$th qubit.
We are interested in the regime where, after a sufficiently large number of time steps $n_{d}$, the system approaches a steady state, i.e., $\rho_{n_{d}}\approx\rho_{n_{d}+1}\approx\rho_{\infty}$. However, we stress that the existence of $\rho_{\infty}$ is not guaranteed for general measurement channels. Nevertheless, for the examples considered in this work, such a steady state indeed exists, as we will show in Secs. IV and VI.
The dynamical process of MDITE involves a competition between two effects: the ITE operator $e^{-\frac{\tau}{2}H}$, which drives the system toward its ground state, and the measurement channel $\mathcal{E}_{p}$, which introduces decoherence and tends to produce a mixed state. As the computational basis is generally not the eigenbasis of $H$, ITE induces quantum fluctuations and reduces classical uncertainty. This suggests potential mixed-state phase transitions with respect to the steady state $\rho_{\infty}$ when tuning the parameters in the protocol.
It is important to note that the time length $\tau$ of each ITE and the measurement rate $p$ serve as primary parameters controlling the behaviors of the steady state $\rho_{\infty}$, as we will discuss in Sec. V. Moreover, Hamiltonian parameters (e.g., coupling strength) also play a crucial role, resulting in a high-dimensional parameter space. This enriches the landscape of behaviors of the output state, rendering the setup of MDITE particularly compelling for investigating and discovering novel many-body physics in open quantum systems.
III Diagrammatic representation of the monitored imaginary-time evolution
Next, we introduce a diagrammatic representation to describe the evolving state $\rho_{n_{d}}$ in the MDITE process. This representation is powerful as it both enhances conceptual understanding and facilitates numerical studies, such as the efficient QMC method we introduce for the MDITE problem in Sec. III.5. Moreover, we note that analytical treatments of the MDITE steady state for general $\tau$ and $p$ are challenging, and only certain continuous limits are tractable, as we will discuss in Sec. V.
[…]
We remark that this section is primarily technical. Readers mainly interested in the results and analysis may skip directly to Sec. IV without loss of continuity.
III.1 ITE without measurements
We begin by examining the case where no measurements are applied throughout the ITE processes, namely when the measurement rate is set to $p=0$. The state after the first time step is
$\rho_{1}\propto e^{-\frac{\tau}{2}H}I_{2^{N}}e^{-\frac{\tau}{2}H}=e^{-\tau H},$ (4)
which can be written as
$|\rho_{1}=\sum_{s,s^{\prime}}\langle s^{\prime}|\rho_{1}|s\rangle\ket{s^{\prime}}\bra{s}$ (5)
[…]
Similarly, the output state after the second step, $\rho_{2}\propto e^{-2\tau H}$, is depicted in Fig. 1(b). We note that the intermediate states $|\ket{r^{\prime}}\bra{r}$ in the representation of $\rho_{2}$ [Fig. 1(b)] correspond to $|\ket{s^{\prime}}\bra{s}$ in the representation of $\rho_{1}$ [Fig. 1(a)], respectively. This follows from $\rho_{2}\propto e^{-\tau/2}\rho_{1}e^{-\tau/2}$, where the two ITE operators act on the upper and lower sides of $\rho_{1}$ in Fig. 1(a). In a manner analogous to tensor-network contraction, these operators propagate $\rho_{1}$, represented by $|\ket{r^{\prime}}\bra{r}$, into $\rho_{2}$, represented by $|\ket{s^{\prime}}\bra{s}$.
Notably, Fig. 1(a) and (b) can be interpreted as two ensembles, due to the summation over all possible $|\ket{s^{\prime}}\bra{s}$ in Eq. (5). The cases of $\rho_{k}$ with $k>2$ follow in complete analogy, with their diagrammatic representations obtained recursively in the same way.
[…]
III.2 ITE with deterministic measurements
We next consider the case in which all qubits are measured deterministically in each time step. Note that the output state $\rho_{1}\propto e^{-\tau H}$ after the first time step is independent of the choice of measurement channel $\mathcal{E}_{p}$ and whether $\mathcal{E}_{p}$ is applied, since measurements on a maximally mixed state always leave it unchanged. Therefore, the diagrammatic representation of $\rho_{1}$ under deterministic measurements remains Fig. 1(a).
In the second time step, we first obtain the measurement-averaged state after the measurement channel $\mathcal{E}_{p}$, which is
$|.\bar{\rho}_{1}=\sum_{s}\langle s|\rho_{1}|s\rangle\ket{s}\bra{s}$. (6)
[…]
After the ITE, the resulting state is
$\rho_{2}\propto e^{-\tfrac{\tau}{2}H}\,\bar{\rho}_{1}\,e^{-\tfrac{\tau}{2}H},$ (7)
which is the output state after the second time step. Analogous to the discussion in Sec. III.1, $\rho_{2}$ can be represented diagrammatically as in Fig. 2(b). Here, the state $|\ket{r}\bra{r}$ corresponds to the state $|\ket{s}\bra{s}$ in $\bar{\rho}_{1}$ shown in Fig. 2(a). It is straightforward to generalize the discussion to a general $\rho_{k}$ $(k>2)$, and the diagrammatic representations also follow recursively.
Figure 2: Diagrammatic representations of the MDITE under deterministic projective measurements. (a) Compared with Fig. 1, the measurement on $\rho_{1}$ identifies the bra and ket components of $\rho_{1}$ in Eq. (5), forming a closed loop for the measurement-averaged state $\bar{\rho}_{1}$. (b) The output state $\rho_{2}$ after the second time step.
III.3 ITE with deterministic measurements on subsystem $A$ only
As the final preparatory step before discussing the general case of probabilistic measurements on each qubit, we consider the scenario in which all qubits in subsystem $A$ are measured projectively with probability one, while the complementary subsystem $B\equiv\overline{A}$ remains unmeasured.
[…]
$|,\bar{\rho}_{1}=\sum_{s_{A},s_{B},s_{B}^{\prime}}\langle s_{A},s_{B}^{\prime}|\rho_{1}|s_{A},s_{B}\rangle\ket{s_{A},s_{B}^{\prime}}\bra{s_{A},s_{B}}$, (8)
where $s_{A}\in\{0,1\}^{\otimes N_{A}}$ and $s_{B}\in\{0,1\}^{\otimes N_{B}}$, with $N_{A}$ and $N_{B}$ denoting the number of qubits in subsystems $A$ and $B$, respectively.
[…]
Figure 3: Diagrammatic representation of the MDITE where only subsystem AA is projectively measured in a deterministic manner. The label “AA” (“BB”) indicates that the diagram applies to all qubits in subsystem AA ($B$). Since the computational basis is local, the evolutions of the reduced density matrices $(\rho_{0})_{A}=\text{Tr}_{B}(\rho_{0})$ and $(\rho_{0})_{B}=\text{Tr}_{A}(\rho_{0})$ can be depicted separately in (a) and (b), respectively.
III.4 ITE with probabilistic measurements
Now we extend the discussions above to a general case in which each site is independently measured with probability $p$ in each time step.
For a system of $N$ qubits and $k$ time steps, the binary choice of measuring or not measuring each qubit in each layer yields $2^{(k-1)N}$ distinct ensembles, which together form an extended ensemble. In this sense, the examples in Figs. 1(b), 2(b), and 3 correspond to three of the $2^{N}$ possible ensembles when $k=2$.
The diagrammatic representation with probabilistic measurements now consists of $N$ components, a natural generalization of Fig. 3, one for each qubit. In Fig. 4, we present an additional example for $N=2$ and $k=3$, illustrating two of the $2^{4}$ possible ensembles.
[…]
III.5 QMC simulations
Importantly, the diagrammatic representation of the evolving state $\rho_{k}$ naturally facilitates QMC studies of $\rho_{n_{d}}$ based on imaginary-time path integrals (after applying $n_{d}$ layers, we obtain the final output state $\rho_{n_{d}}$), such as the stochastic series expansion (SSE) [77, 79, 81, 93, 82, 116, 118], as long as the Hamiltonian is sign-problem-free. In addition, it also makes it straightforward to implement tensor-network–based simulations [75].
[…]
Therefore, we have
$\mathcal{Q}_{n_{d}}=\sum_{\{A_{k}\}}\prod_{l=2}^{n_{d}}\bigg[p^{N_{A_{l}}}(1-p)^{N-N_{A_{l}}}\bigg]Q_{n_{d}}(\{A_{k}\}),$ (9)
where $N_{A_{l}}$ denotes the number of qubits in subsystem $A_{l}$. Since different ensembles $\{Q_{n_{d}}(\{A_{k}\})\}$ carry different weights determined by the measurement rate $p$, transitions between ensembles can be implemented in QMC simulations by designing appropriate selection probabilities that satisfy the detailed balance condition.
Within the SSE framework, we have developed an unbiased and efficient QMC algorithm to simulate Eq. (9), applicable to Hamiltonians for ITEs in arbitrary spatial dimensions. While the present work focuses on the steady state $\rho_{\infty}$ ($n_{d}\to\infty$), the algorithm also enables studies of the dynamics of the MDITE process. Further details of the QMC algorithm are provided in Appendices A and B.
IV Numerical results for a 1D Model with discrete symmetry
IV.1 Model and observables
As the first example, we consider the one-dimensional (1D) transverse-field Ising model (TFIM) [72] as the Hamiltonian in the MDITE protocol, which is given by
$H=-\sum_{i}Z_{i}Z_{i+1}-h\sum_{i}X_{i},$ (10)
where $Z_{i}$ and $X_{i}$ are the Pauli operators acting on site $i$, and $h$ is the transverse field strength. The computational basis is chosen to be the $Z$-basis. The length of the chain is denoted by $L\equiv N$, and the periodic boundary condition, i.e., $N+1\equiv 1$, is considered.
This model has a global $\mathbb{Z}_{2}$ symmetry and the ground state exhibits a quantum phase transition at $h_{c,\mathrm{GS}}=1$, as determined by the Kramers-Wannier duality, which belongs to the 2D classical Ising universality class. For $h<h_{c,\mathrm{GS}}$, the system is in a ferromagnetic (FM) ordered phase with nonzero magnetization, while for $h>h_{c,\mathrm{GS}}$, it is in a paramagnetic (PM) disordered phase with zero magnetization.
Though free energy is not a well-defined observable in the MDITE protocol, the magnetization of spins, $\langle m\rangle\equiv\langle\sum_{i=1}^{N}Z_{i}/N\rangle$, remains accessible and physical, serving as an important diagnosis for characterizing the steady state $\rho_{\infty}$.
Specifically, for the TFIM, we numerically calculate the absolute magnetization $\langle|m|\rangle$, while also considering the second moment $\langle m^{2}\rangle$ and the fourth moment $\langle m^{4}\rangle$, from which the dimensionless Binder ratio $R_{2}=\langle m^{4}\rangle/\langle m^{2}\rangle^{2}$ is obtained. Note that $\langle|m|\rangle$ does not correspond to $\langle|\sum_{i=1}^{N}Z_{i}/N|\rangle$. Instead, in QMC simulations, $\langle|m|\rangle$ represents the ensemble average of the absolute magnetization measured in individual classical QMC samples [81]. For the standard TFIM, observables $\langle|m|\rangle$ and $R_{2}$ are important for identifying the phase transition point from finite-size simulations [4, 36, 80, 128, 117].
In the FM phase, the magnetization distribution is double-peaked around $\pm\langle|m|\rangle$, serving as a direct signature of spontaneous symmetry breaking (SSB). This behavior leads to $R_{2}\to 1$, while in the PM phase the distribution becomes Gaussian, yielding $R_{2}\to 3$. Remarkably, as we will show below, these observables can also be used to characterize the mixed-state phase transitions of $\rho_{\infty}$ in the presence of measurements.
The current section mainly presents the numerical results obtained from QMC simulations. The corresponding discussions and theoretical analysis will be provided in Sec. V.
IV.2 Steady state
We here demonstrate the existence of the steady state in the MDITE protocol, as noted in Sec. II. Fig. 5 shows that for the control parameters $(\tau,h,p)=(1,1.8,0.66)$—which we will later identify as a critical point—the state $\rho_{n_{d}}$ converges to a steady state in the limit $n_{d}\to\infty$.
This is evidenced by the rapid convergence of both $\langle m^{2}\rangle$ and the Binder ratio $R_{2}$ with increasing the number of time steps, producing dynamics reminiscent of the quantum Zeno effect [62, 39, 42]. We observe similar convergence behavior for other parameter choices.
In practice, we find that $n_{d}=2L/\tau$ is sufficient to reach the steady state for the 1D TFIM, which implies that the dynamical exponents for the mixed-state critical points satisfy that $z\leq 1$. As we will show in Sec. IV.5, our numerical simulations give $z\approx 1$ for all the critical points investigated here.
Figure 5: When setting $(\tau,h,p)=(1,1.8,0.66)$ for the MDITE with the 1D TFIM: (a) Convergence of the second moment $\langle m^{2}\rangle$ with increasing the number of time steps for various system sizes $L$. (b) Convergence of the Binder ratio $R_{2}$ with increasing time steps for various system sizes $L$.
IV.3 Mixed-state MIPT
[…]
Figure 6: For $(\tau,h)=(1,1.8)$ in the MDITE of the 1D TFIM: (a) Binder ratios $R_{2}$ for different system sizes $L$ cross at $p_{c}\approx 0.667$; (b) The inset panel shows $\langle|m|\rangle$ as a function of $p$ for various $L$, while the main panel presents the finite-size scaling and data collapse of $\langle|m|\rangle$.
For $(h,p)=(2.5,0.5)$: (c) Binder ratios $R_{2}$ for different system sizes $L$ cross at $\tau_{c}\approx 0.265$; (d) $\langle|m|\rangle$ as a function of $\tau$ for various $L$ and data collapse of $\langle|m|\rangle$.
For $(\tau,p)=(1.2,0.8)$, (e) Binder ratios $R_{2}$ for different system sizes $L$ cross at $h_{c}\approx 1.84$; (f) $\langle|m|\rangle$ as a function of $h$ for various $L$ and data collapse of $\langle|m|\rangle$.
We first explore the three-dimensional parameter space $(\tau,h,p)$ by fixing $\tau$ and $h$, and then scanning the measurement rate $p$ to examine whether it induces a phase transition. Our extensive numerical calculations show that tuning $p$ indeed drives a transition for a wide range of $(\tau,h)$ values when $h>1$.
One example is shown for $(\tau,h)=(1,1.8)$ in Fig. 6(a), where a clear crossing of the Binder ratio $R_{2}$ for different system sizes $L$ identifies a critical point $p_{c}$. The asymptotic behavior of $R_{2}$—approaching $3$ for $p<p_{c}$ and $1$ for $p>p_{c}$—provides strong evidence for a mixed-state phase transition. Moreover, since the measurement channel preserves the weak (or average) $\mathbb{Z}_{2}$ symmetry of the state [76], the observed transition supported by the behavior of binder ratios thus corresponds to a weak-to-nothing SSB, analogous to thermal phase transitions in the 2D classical Ising model. This separates a disordered mixed phase for $p<p_{c}$ and an ordered mixed phase for $p>p_{c}$.
Note that when $p=0$, the steady state reduces to the ground state of the 1D TFIM, which is in the PM phase when $h=1.8>1$. The results in Fig. 6(a) thus demonstrate that sufficiently strong projective measurements can drive the system into a mixed FM phase, polarized along the $z$-axis. In contrast to measurement-induced entanglement-entropy transitions that require quantum trajectory analysis [89, 6], this transition is detectable using conventional linear observables without the need for information-theoretic diagnostics.
To identify the universality class of the critical point, we perform a finite-size scaling analysis [74] on
$x_{L}=\frac{p-p_{c}}{p_{c}}L^{1/\nu},\quad y_{L}=\langle|m|\rangle L^{\beta/\nu},$ (11)
where $\nu$ and $\beta$ denote the critical exponents related to the correlation length and the magnetization. From the data collapse, we extract $p_{c}\approx 0.667$, $\nu\approx 1.08$, and $\beta\approx 0.43$, whose numerical stability is reported in Table 1. The corresponding plots are presented in Fig. 6(b).
Importantly, the critical exponents $\nu$ and $\beta$ are quite different from that of the equilibrium 2D or 3D classical Ising criticality, which means this mixed-state phase transition is not a standard Ising phase transition, though the numerical results indicate an Ising-like $\mathbb{Z}_{2}$ SSB.
Parameters $L_{\min}$ Critical point $\nu$ $\beta$
$(\tau,h)=(1,1.8)$ 112 $p_{c}=0.6668(6)$ 1.073(5) 0.422(8)
128 $p_{c}=0.6667(7)$ 1.077(6) 0.424(9)
144 $p_{c}=0.6665(9)$ 1.079(7) 0.43(1)
160 $p_{c}=0.666(1)$ 1.084(8) 0.43(1)
$(h,p)=(2.5,0.5)$ 112 $\tau_{c}=0.2651(1)$ 1.1873(2) 0.480(1)
128 $\tau_{c}=0.2652(1)$ 1.189(1) 0.480(2)
144 $\tau_{c}=0.2651(2)$ 1.190(2) 0.480(4)
160 $\tau_{c}=0.2652(4)$ 1.191(6) 0.480(7)
$(\tau,p)=(1.2,0.8)$ 112 $h_{c}=1.8375(3)$ 1.181(5) 0.46(1)
128 $h_{c}=1.8373(7)$ 1.18(1) 0.46(1)
144 $h_{c}=1.837(1)$ 1.19(2) 0.46(1)
160 $h_{c}=1.835(2)$ 1.21(3) 0.46(1)
Table 1: Critical exponents obtained from data collapse of $\langle|m|\rangle$ with $L=112,128,144,160,176,192$. To test the stability of the extracted exponents, we first discard the $L=112$ data, setting $L_{\min}=128$, and then progressively increase $L_{\min}$ to 160160 to retain only the three largest system sizes.
IV.4 $\tau$- and $h$-induced phase transitions
In addition to the MIPTs driven by tuning $p$, the steady state exhibits mixed-state phase transitions when either the imaginary-time step $\tau$ or the field strength $h$ is varied, with the other parameters fixed.
As an example, for $(h,p)=(2.5,0.5)$, the Binder ratio $R_{2}$ shows a clear crossing at $\tau_{c}\approx 0.265$, with a scaling analysis of $\langle|m|\rangle$ yielding $\nu\approx 1.19$, $\beta\approx 0.48$ [Fig. 6(c, d), Table 1].
Similarly, for $(\tau,p)=(1.2,0.8)$, we identify a transition at $h_{c}\approx 1.84$, characterized by $\nu\approx 1.18$ and $\beta\approx 0.46$ [Fig. 6(e, f), Table 1]. Notably, the values of $\nu$ and $\beta$ change for different critical points.
IV.5 Dynamical scaling
To better understand these mixed-state phase transitions, we also extract the dynamical exponent $z$ by simulating the relaxation dynamics from the initial state to the steady state. Specifically, we determine the relaxation time $\xi_{\tau}$ by fitting the convergence of the order parameter $\langle|m_{z}|\rangle$ as a function of the number of time steps $n_{d}$ using the fitting formula
$m(n_{d})=a\exp(-n_{d}/\xi_{\tau})+b,$ (12)
for each system size. At the critical point, the relaxation time is expected to scale as $\xi_{\tau}\propto L^{z}$.
For the measurement-induced critical point $(\tau,h,p)=(1,1.8,0.6667)$ in Sec. IV.3, we obtain $z\approx 1$. The data are shown in Fig. 7, and the stability analysis yields consistent results, as summarized in Table 2.
Figure 7: When setting $(\tau,h,p)=(1,1.8,0.6667)$ for the MDITE with the 1D TFIM: (a) The relaxation process of $\langle|m|\rangle$ with increasing the number of time steps $n_{d}$ for various system sizes $L$. It satisfies the formula $m=a\exp(-n_{d}/\xi_{\tau})+b$, where $a$, $\xi_{\tau}$ and $b$ are fitting parameters. (b) Extracting the dynamical exponent $z$ from $\xi_{\tau}$ with various system size $L$. The fitting formula is $\xi^{-1}_{\tau}=kL^{-z}$, where $k$ and $z$ are fitting parameters.
$L_{\min}$ $k$ $z$
64 7.1(2) 0.997(5)
72 6.9(2) 0.993(6)
80 6.7(3) 0.987(7)
88 6.5(3) 0.978(9)
96 6.8(4) 0.99(1)
Table 2: For $(\tau,h,p)=(1,1.8,0.6667)$, we extract the dynamical exponent $z$ from the relaxation correlation length using system sizes L=64,72,,128L=64,72,...,128. To test the stability of the extracted exponent, small system sizes are gradually discarded. $L_{\min}$ denotes the smallest systems size used in each fitting.
Similarly, we extract the dynamical exponents for the other two parameter sets $(\tau,h,p)=(1.2,1.8373,0.8)$ and $(0.4675,2.5,0.5)$ in Sec. IV.4, where the phase transitions are driven by the transverse field $h$ and the imaginary-time step $\tau$, respectively. As reported in Appendix F, the exponents are also close to one.
After accounting for finite-size effects, our numerical results are consistent with a dynamical critical exponent $z\approx$1, suggesting an emergent space–time isotropy at the steady-state critical points. This behavior is compatible with an emergent Lorentz-invariant description and possibly an effective conformal field theory, although further evidence is required to substantiate this interpretation.
V Analysis of the criticalities in MDITE steady states
Let us now analyze the novel mixed-state phase transitions observed in the numerical simulations presented in the previous section.
Note that there are three tractable continuous limits: the continuous-time limit ($\tau\to 0$, $p\sim O(1)$), the weak-measurement limit ($p\to 0$, $\tau\sim O(1)$), and the combined limit ($\tau\to 0$, $p\to 0$, with $\tau\sim p$). Away from these limits, analytical description could be challenging.
As we show below, all three limits give rise only to the conventional Ising universality class. Particularly, neither the continuous-time limit nor the weak-measurement limit yields results that are even qualitatively consistent with the numerical findings obtained away from these regimes. For this reason, we defer detailed perturbative calculations and discussions of these two limits to Appendices C and D, respectively.
By contrast, the combined limit provides qualitatively consistent behavior, although the resulting universality class remains that of the standard Ising transition. We thus conclude that, sufficiently far from these continuous limits (such as the numerical results in Sec. IV), the mixed-state phase transitions may fall outside the scope of known universality classes.
V.1 Choi representation
We first represent the steady state in the double-Hilbert space formalism. Using the identities
[…]
the local measurement channel in Eq. (3) can be rewritten as
$\mathcal{E}_{p,i}(\rho)=\bigg(1-\frac{p}{2}\bigg)\rho+\frac{p}{2}Z_{i}\rho Z_{i}.$ (14)
Under the Choi-Jamiołkowski isomorphism [13, 40], $\mathcal{E}_{p,i}$ can be mapped onto a superoperator acting on a doubled Hilbert space, which is
$\mathcal{E}_{p,i}\longrightarrow\hat{\mathcal{E}}_{p,i}=(1-p)^{1/2}e^{\tilde{\tau}\,Z_{i}^{L}Z_{i}^{R}},$ (15)
where the inter-copy coupling strength is given by
$\tilde{\tau}\equiv\tilde{\tau}(p)=\tanh^{-1}\!\left(\frac{p}{2-p}\right),$ (16)
and superscripts $L$ and $R$ denote operators acting on the left and right copies of the Hilbert space, respectively. In addition, for $x\equiv p/(2-p)\in[0,1]$, we can rewrite
$\tilde{\tau}=\tanh^{-1}(x)=\frac{1}{2}\ln\bigg(\frac{1+x}{1-x}\bigg)=-\frac{1}{2}\ln(1-p).$ (17)
When $\tilde{\tau}\to 0$, $\tilde{\tau}$ can be approximately replaced by $p/2$.
Accordingly, the evolution described in Sec. II becomes
$\left|\rho_{k}\right\rangle\!\rangle\propto e^{-\tau\tilde{H}}e^{\tilde{\tau}\sum_{i}Z_{i}^{L}Z_{i}^{R}}\left|\rho_{k-1}\right\rangle\!\rangle,$ (18)
with the two-copy Hamiltonian defined as
$\tilde{H}\equiv\frac{1}{2}(H^{L}+H^{R}).$ (19)
V.2 Effective Hamiltonian in the combined limit
In the limit $\tau\to 0$ and $\tilde{\tau}\to 0$, with $\tau\sim\tilde{\tau}$, Eq. (18) can be simplified using the Baker-Campbell-Hausdorff (BCH) formula. Keeping terms up to linear order in $\tau$ and neglecting $O(\tau^{2})$ corrections, we obtain
$e^{-\tau\tilde{H}}e^{\tilde{\tau}\sum_{i}Z_{i}^{L}Z_{i}^{R}}\approx e^{-\tau H_{\text{eff}}},$ (20)
where the effective Hamiltonian is given by
$H_{\text{eff}}\equiv\tilde{H}-\frac{p}{2\tau}\sum_{i}Z_{i}^{L}Z_{i}^{R}.$ (21)
Therefore, the ITE and measurement operations can be combined into a single effective transformation in this limit, which is exactly an doubled-space ITE with time length $\tau$, under an effective two-copy Hamiltonian $H_{\mathrm{eff}}$.
Moreover, the initial state in the doubled space is given by
[…]
For the example of the 1D TFIM (10), the effective Hamiltonian takes the form
${H}_{\text{eff}}=$ $-\frac{1}{2}\sum_{\langle ij\rangle}(Z_{i}^{L}Z_{j}^{L}+Z_{i}^{R}Z_{j}^{R})-\frac{h}{2}\sum_{i}(X_{i}^{L}+X_{i}^{R})$ (23)
$-\frac{p}{2\tau}\sum_{i}Z_{i}^{L}Z_{i}^{R}.$
When $p/\tau=0$, the two copies are decoupled and the $H_{\text{eff}}$ has a $\mathbb{Z}_{2}^{L}\times\mathbb{Z}_{2}^{R}$ symmetry, where $\mathbb{Z}_{2}^{L(R)}$ corresponds to independent spin-flip symmetry in the left (right) copy, generated by $X^{L(R)}\equiv\prod_{i}X_{i}^{L(R)}$.
In the combined limit $\tau\sim\tilde{\tau}\sim p$, the inter-copy coupling is finite, and only the diagonal spin-flip transformation preserves the Hamiltonian. The symmetry is therefore reduced to the diagonal subgroup $\mathbb{Z}_{2}^{\text{diag}}=\{1,X^{L}X^{R}\}\cong\mathbb{Z}_{2}$, and the effective Hamiltonian (23) is equivalent to a standard TFIM defined on a two-leg ladder geometry, as illustrated in Fig. 8(a).
Figure 8: (a) Ladder geometry of $H_{\text{eff}}$ in Eq. (23) for the case where $H$ is the 1D TFIM. The two legs correspond to the left ($L$) and right ($R$) copies of the chain (two rings under periodic boundary conditions). The interlayer Ising couplings along the rungs have strength $p/2\tau$. (b) Ground-state phase diagram of $H_{\text{eff}}$ obtained from QMC simulations.
Fig. 8(b) shows the ground-state phase diagram of $H_{\text{eff}}$ in Eq. (23). The critical point at $h=1$ lies at $p/\tau=0$. Note that the combined limit requires $p/\tau\sim O(1)$, and this corresponds to the region $h>1$. Thus, sufficiently large $p/\tau$ can drive a transition from the disordered phase to the ordered FM phase when $h>1$ in the doubled space.
V.3 MIPT in the combined limit
Note that in the effective Hamiltonian (23), both the interlayer Ising coupling (from measurements) and the intralayer Ising coupling (from the Ising interaction in the original Hamiltonian) favor FM ordering in the same diagonal $\mathbb{Z}_{2}$ channel, while the transverse field $h$ enhances quantum fluctuations and tends to disorder the system.
The key to understanding the mixed-state MIPT in the combined limit is the dual role of the $Z$-basis: it is both the basis in which projective measurements are performed and the basis in which the ordered phase of the Hamiltonian is defined. When $p/\tau$ is sufficiently large compared to hh—equivalently, when the decoherence strength $p$ dominates over the coherence scale $\tau h$ via the ITE operator $e^{-\frac{\tau}{2}H}$—the system develops FM long-range order for the steady state in the doubled space.
This ordered phase is characterized by the local order parameter $\langle\!\langle\rho_{\infty}|Z_{i}^{L(R)}|\rho_{\infty}\rangle\!\rangle$ (or one can equivalently consider the magnetization averaged over all sites), where $\left|\rho_{\infty}\right\rangle\!\rangle$ denotes the ground state of $H_{\text{eff}}$. In the original Hilbert space, the Choi mapping relates this quantity to the nonlinear quantity $\mathrm{Tr}(\rho_{\infty}^{2}Z_{i})$, with $\rho_{\infty}$ denoting the steady state of the MDITE. Though the parameter region where $\mathrm{Tr}(\rho_{\infty}^{2}Z_{i})\neq 0$ needs not coincide exactly with that where the linear order parameter $\mathrm{Tr}(\rho_{\infty}Z_{i})$ is nonzero, both capture the SSB physics. The resulting mixed-state MIPT for the steady state of the MDITE protocol in the combined limit is therefore expected to be qualitatively identical and within the same 2D classical Ising universality class.
We emphasize that the results in the combined limit are qualitatively consistent with the numerical results in Sec. IV.3 away from the continuous limits. The mechanism of these mixed-state phase transitions in Sec. IV.3 could therefore be understood as follows.
V.4 MIPT away from the continuous limits
In the regime $h>1$ with relatively short imaginary-time length ($\tau h\lesssim p$), frequent projective measurements at large $p$ interrupt the buildup of quantum coherence. The transverse field, which generates quantum fluctuations in the $X$ basis, is repeatedly suppressed before correlations can develop, as the state is frequently projected onto the $Z$ basis. With off-diagonal coherence effectively suppressed, the dynamics become dominated by the diagonal Ising terms, steering the system toward a FM phase analogous to that of the classical Ising model.
In contrast, when the transverse field is sufficiently strong or the evolution time is sufficiently long ($\tau h\gtrsim p$), quantum coherence develops rapidly within each imaginary-time interval. Coherence is established faster than it can be suppressed by measurements, preventing the formation of FM order even at $p=1$. Consequently, no phase transition occurs in this regime upon tuning the measurement rate $p$. One may also expect that when $h\leq 1$, no phase transition characterized by spin magnetization occurs upon tuning $p$, which has also been observed in our numerical observations.
Since the competition between quantum coherence and measurement-induced decoherence can establishe a critical balance, tuning the system parameters, including $p$, $\tau$, and $h$, toward this balance drives a mixed-state phase transition separating the ordered and disordered phases, as shown in Sec. IV.4 In this sense, transitions induced by varying $p$, $\tau$, or $h$ originate from the same underlying mechanism. The difference lies only in which parameter controls the competition.
Furthermore, while the combined limit admits a description in terms of the effective Hamiltonian $H_{\mathrm{eff}}$ and exhibits critical behavior consistent with the Ising universality class, the numerical results in the intermediate regime (away from the three limits) show critical exponents that do not match those of any known universality class. In particular, the extracted exponents appear to vary continuously as the parameters $h$, $\tau$, and $p$ are changed. These observations suggest the possibility that the mixed-state phase transitions in the intermediate regime belong to some distinct universality class that has not been previously identified.
Physically, the nontrivial mixed-state phase transitions should arise from the intrinsically nonequilibrium nature of the problem. Based on the BCH expansion, one can also define an effective Hamiltonian $H_{\text{eff}}$ through
$e^{-\tau\tilde{H}}e^{\tilde{\tau}\sum_{i}Z_{i}^{L}Z_{i}^{R}}\equiv e^{-\tau H_{\text{eff}}}.$ (24)
However, In general, the resulting effective Hamiltonian $H_{\text{eff}}$ could become highly nonlocal due to the infinite series of nested commutators, resulting critical behavior therefore fall outside the standard equilibrium classification.
In addition, the emergence of long-range order with an increasing measurement rate can also be understood intuitively from the perspective of QMC simulations: a higher measurement rate naturally generates larger clusters that connect a broader range of sites in each QMC update. This provides a direct and intuitive picture of how measurements drive ordering in the system of TFIM, highlighting the use of the diagrammatic representations in Sec. III and the advantage of QMC in revealing the underlying mechanism. Further details are provided in Appendix. E.
Figure 9: Away from the continuous limits, critical surfaces of the steady state of the MDITE protocol viewed from different angles [(a) and (b)]. To facilitate the distinction of critical $p$ corresponding to different parameters on the critical surface, the figure employs a colorbar to map $p$ to different colors.
V.5 Critical surface away from the continuous limits
To visualize the phase transitions of the MDITE steady states, we construct a schematic phase diagram based on approximate transition points obtained from Binder ratio crossings in several system sizes, as shown in Fig. 9. Importantly, Fig. 9(b) shows that the critical surface shifts toward larger $\tau$ as $h$ approaches $1$. This trend is expected: longer evolution times require weaker transverse fields to allow FM order to emerge under strong measurements. Consequently, in the limit $h\to 1$, the critical surface is expected to extend infinitely along the $\tau$-axis, permitting a transition at any $\tau$ by tuning $p$. A similar reasoning applies in the opposite limit, $h\to\infty$ and $\tau\to 0$.
VI Two-dimensional model with continuous symmetry
VI.1 Model and observables
As the second example, we consider the 2D $S=1/2$ columnar dimerized Heisenberg model (CDHM) on a $L\times L$ square lattice [59]. The Hamiltonian is given by
$H=\sum_{\langle ij\rangle_{1}}\mathbf{S}_{i}\cdot\mathbf{S}_{j}+g\sum_{\langle ij\rangle_{g}}\mathbf{S}_{i}\cdot\mathbf{S}_{j}$ (25)
where $\mathbf{S}_{i}$ is the spin operator at site $i$, and $\langle ij\rangle_{1}$ and $\langle ij\rangle_{g}$ denote two different sets of nearest-neighbor pairs, as shown in Fig. 10. We also consider the local $Z$-basis as the computational basis, and the periodic boundary condition is chosen. The model has a global $SU(2)$ symmetry and the ground state exhibits a quantum phase transition at $g_{c,\mathrm{GS}}\approx 1.90951$ [81, 59, 55], which belongs to the 3D $O(3)$ universality class. The critical point separates a Néel ordered phase ($g<g_{c,\mathrm{GS}}$) from a dimerized phase ($g>g_{c,\mathrm{GS}}$).
Figure 10: Square lattice of the 2D CDHM, where the red thick bonds denote couplings with strength $g$, while the remaining bonds have unit strength.
Due to the antiferromagnetic (AFM) nature of the Heisenberg interactions, we consider the staggered magnetization $\langle m_{z}\rangle=\Big\langle\sum_{i}(-1)^{x_{i}+y_{i}}Z_{i}/N\Big\rangle$ as the order parameter, where $x_{i}$ and $y_{i}$ denote the integer coordinates of site $i$ in the lattice. As in the TFIM, we compute the absolute magnetization $\langle|m_{z}|\rangle$, the second and fourth moments of the magnetization, and the Binder ratio $R_{2,z}$. Furthermore, for the ground state of the CDHM, the Binder ratio associated with the $O(3)$ phase transition approaches $R_{2,z}\to 9/5$ in the Néel phase and $R_{2,z}\to 3$ in the dimerized phase [78, 105, 17, 55].
VI.2 Steady state
As in the 1D TFIM example, the MDITE protocol in this case also drives the system toward a steady state in the limit of large number of time steps $n_{d}$. Fig. 11 illustrates this convergence for representative parameters $(\tau,g,p)=(1,3.5,0.5)$, where both the staggered magnetization $\langle|m_{z}|\rangle$ and the Binder ratio $R_{2,z}$ quickly approach their stationary values. In practice, we again find that $n_{d}=2L/\tau$ is sufficient to reach the steady state, and this choice is used in all subsequent QMC simulations for the CDHM. As discussed later, the dynamical exponent $z$ for the mixed-state phase transitions in this example is also around one.
Figure 11: When setting $(\tau,g,p)=(1,3.5,0.5)$ for the MDITE with the 2D CDHM: (a) Convergence of the squared magnetization $\langle m^{2}\rangle$ with increasing the number of time steps $n_{d}$ for various system sizes $L$. (b) Convergence of the Binder ratio $R_{2,z}$ with increasing $n_{d}$ for various system sizes $L$.
VI.3 Mixed-state phase transitions
[…]
Figure 12: For $(\tau,g)=(1,3.5)$ in the MITE of the 2D CDHM: (a) Binder ratios $R_{2,z}$ for different system sizes $L$ cross at $p_{c}\approx 0.354$. (b) The inset panel shows $\langle|m_{z}|\rangle$ as a function of $p$ for various $L$, while the main panel presents the finite-size scaling and data collapse of $\langle|m_{z}|\rangle$.
For $(g,p)=(3,0.1)$: (c) Binder ratios $R_{2,z}$ for different system sizes $L$ cross at $\tau_{c}\approx 0.468$; (d) $\langle|m|\rangle$ as a function of $\tau$ for various $L$ and data collapse of $\langle|m|\rangle$.
For $(\tau,p)=(2,0.3)$: (e) Binder ratios $R_{2,z}$ for different system sizes $L$ cross at $g_{c}\approx 2.8$; (f) $\langle|m|\rangle$ as a function of $g$ for various $L$ and data collapse of $\langle|m|\rangle$.
(g) and (h): Critical surfaces viewed from different angles.
The parameter space of the steady state in this model is spanned by $(\tau,g,p)$. Compared with the TFIM Hamiltonian, the CDHM Hamiltonian not only provides a 2D example but also exhibits a global continuous $SU(2)$ symmetry. Interestingly, even at $g=0$, the ITE operator contains off-diagonal components $(S^{+}_{i}S^{-}_{j}+S^{-}_{i}S^{+}_{j})/2$. Increasing $g$ introduces both diagonal and off-diagonal terms, in sharp contrast to the TFIM case where tuning the Hamiltonian parameter $h$ only modifies the off-diagonal part. Following the same strategy as in the TFIM case, we fix two of the parameters in $\{\tau,g,p\}$ and vary the third to search for mixed-state phase transitions.
Taking the example of tuning $p$ at fixed $(\tau,g)=(1,3.5)$ in Fig. 12(a) and (b), we observe a clear crossing of the Binder ratio $R_{2,z}$ at $p_{c}\approx 0.354$, with $R_{2,z}\to 3$ for $p<p_{c}$ and $R_{2,z}\to 1$ for $p>p_{c}$. If $p=0$, the system is in the dimerized phase which has no order in the staggered magnetization. This indicates a mixed-state phase transition from a disordered phase to an AFM Ising-like ordered phase along the $z$-axis, driven by increasing the measurement strength.
Although the Binder ratio exhibits Ising-like characteristic, similar to the case of TFIM, the extracted critical exponents in the case of 2D CDHM again do not correspond to any known universality class including the 3D Ising universality class for a closed system, as shown in Table 3. Similar phenomena can also be observed when tuning $g$ and $\tau$, as shown in Fig. 12(c-f) and Table 3. Importantly, the critical exponents $\nu$ and $\beta$ also change as we tune $p$, $\tau$, or $g$.
Parameters $L_{\min}$ Critical point $\nu$ $\beta$
$(\tau,g)=(3.5,1)$ 24 $p_{c}=0.3537(4)$ 0.714(7) 0.650(11)
28 $p_{c}=0.3537(4)$ 0.716(6) 0.651(11)
32 $p_{c}=0.3538(5)$ 0.717(6) 0.650(12)
$(g,p)=(3,0.1)$ 24 $\tau_{c}=0.4679(5)$ 0.799(6) 0.718(9)
28 $\tau_{c}=0.4679(6)$ 0.801(6) 0.721(10)
32 $\tau_{c}=0.4678(7)$ 0.801(6) 0.720(12)
$(\tau,p)=(2,0.3)$ 24 $g_{c}=2.806(3)$ 0.736(18) 0.654(26)
28 $g_{c}=2.806(2)$ 0.739(17) 0.653(26)
32 $g_{c}=2.806(2)$ 0.741(15) 0.650(26)
Table 3: Finite-size stability analysis of the critical exponents obtained from data collapse of $\langle|m|\rangle$ in the CDHM. System sizes are $L=24,28,32,40,48$. Small system sizes are progressively discarded to test the stability of the extracted critical exponents.
We also extract dynamical exponents $z\approx 1$ for all three quantum critical points. These results are consistent with approximately isotropic space-time scaling and are compatible with an emergent Lorentz-invariant description. Further details are provided in Appendix G.
VI.4 Analysis of the phase transitions
To further interpret these results, note that any nonzero measurement rate $p$ explicitly breaks the (weak) $SU(2)$ symmetry to a residual (weak) $U(1)\rtimes\mathbb{Z}_{2}$ symmetry, favoring the $z$-axis.
Although increasing $g$ in the Hamiltonian introduces both diagonal and off-diagonal terms, only the diagonal components are compatible with the projective measurements along the $z$-axis, which correspond to the AFM Ising interactions. In Appendix. H, we also present the results demonstrating the absence of off-diagonal long-range order in this model.
To understand this kind of phase transitions, one can similarly construct an effective two-copy Hamiltonian in the combined limit $\tau\to 0$, $p\to 0$, with $\tau\sim p$, as that in Sec. V. In this limit, the resulting effective model corresponds to a bilayer geometry consisting of two CDHM layers coupled by interlayer Ising interactions. When each individual layer is in the dimerized disordered phase ($g>g_{c,\text{GS}}$), tuning the measurement rate $p$ (equivalently, the interlayer coupling strength) can drive Ising phase transitions associated with the 3D classical Ising universality class.
Away from the continuous limits, when $g>g_{c,\mathrm{GS}}$, and the imaginary-time length is short ($\tau g\lesssim p$), a high measurement rate $p$ suppresses the off-diagonal coherence generated by the ITE operator, leading to dynamics dominated by the diagonal AFM Ising terms. This drives the system into an ordered phase with nonzero staggered magnetization along the $z$-axis. In contrast, if $\tau g\gtrsim p$, the off-diagonal coherence builds up rapidly, and the final steady state remains disordered even at $p=1$.
Therefore, for projective measurements along the $z$-axis, the MDITE with the 2D CDHM generalizes the physics of the 1D TFIM discussed in Secs. IV and V to 2D. The associated phase transitions also appear to lie beyond conventional universality classes.
Since the CDHM can also be simulated using cluster updates within QMC, the emergence of long-range orders induced by measurements can also be understood through the cluster formation in Appendix. E. We emphasize that it would be interesting to explore other types of decoherence channels that preserve other symmetries, such as the $SU(2)$ symmetry, which we leave for future work. In such cases, more advanced update algorithms beyond conventional cluster methods for QMC simulations would be necessary, potentially revealing qualitatively different types of phase transitions.
The schematic critical surface associated with this model is shown in Fig. 12(g) and (h). This critical surface also shows that large $g$ requires small evolution time, while long evolution time requires weak coupling $g$. This inverse relationship between $g$ and $\tau$ closely resembles that of the 1D TFIM.
Moreover, we note that since the state still has a $U(1)$ symmetry, off-diagonal long-range order may emerge if the imaginary-time step length in the MDITE protocol is allowed to scale with the system size, i.e., $\tau\propto L^{\alpha}$ for a sufficiently large $\alpha$. In this regime and for $g<g_{c,\text{GS}}$, the final step of the MDITE protocol (measurement followed by ITE) effectively reduces to a conventional ITE procedure for preparing the ground state or low-energy excited state of the Hamiltonian, with the initial state to be $\mathcal{E}_{p}(\rho_{n_{d}-1})$. This regime represents an interesting extension of the present framework and may be explored in future studies.
VII Conclusion and discussions
In this work, we have introduced MDITE as a versatile framework for studying novel mixed-state phases and phase transitions in open quantum systems. By combining ITE with probabilistic projective measurements, MDITE enables the exploration of fully non-unitary dynamics and their steady states, providing both conceptual insight and numerical accessibility for studying high-dimensional open quantum systems.
As illustrative examples, we applied MDITE to the 1D TFIM and the 2D CDHM, uncovering distinct mixed-state phases and associated phase transitions in the steady states of the protocol. Remarkably, the Ising-like critical behavior observed in both 1D and 2D models may, to the best of our knowledge, represent new universality classes of mixed-state phase transitions in open quantum systems. We also developed a theoretical description in the important combined limit in Sec. V, which yields results qualitatively consistent with the numerical observations. In addition, the transitions admit a natural interpretation in terms of cluster formation within the QMC simulations, providing an intuitive picture of the emergent ordering in the mixed states.
At the methodological level, we introduced a diagrammatic representation of the MDITE process, which facilitates QMC and other numerical simulations on a broad class of many-body Hamiltonians, including those exhibiting SSB, topological phase transitions, and topologically ordered phases. While our analysis focused on magnetization and the Binder ratio, the computation of information-theoretic quantities—such as entanglement entropy [35, 38, 61, 15, 127, 18, 108, 96, 114, 46, 57, 107] and Rényi negativity (including entanglement witnesses based on partitial transpose moments) [19, 112, 1, 104, 14, 120, 22, 67, 97]—is straightforward in combination with recent QMC techniques.
Our work opens several promising directions for future research. First, establishing the precise universality classes of the mixed-state phase transitions observed here remains an important open problem. In particular, it would be valuable to understand how these critical points, which appear to go beyond conventional Ising universality, connect to standard Ising critical points in the three continuous limits in Sec. V.
Second, it would be interesting to extend the MDITE framework to more complex Hamiltonians and more general decoherence channels beyond projective measurements, thereby accessing a broader range of many-body phenomena. In particular, long-range entanglement or topological order in mixed states [52, 53, 24, 54, 85, 45], as well as strong-to-weak spontaneous symmetry breaking [43, 44, 76, 29, 37, 109, 33, 31, 32, 124, 91, 50, 126, 83], can be explored within the MDITE framework.
While we have characterized the dynamical critical exponent in the present work, it would be interesting to further explore more detailed dynamical properties of the protocol. In particular, a systematic study of relaxation dynamics and critical scaling toward the steady state may reveal richer structures and additional universal features beyond those captured by the stationary mixed states.
On the experimental side, MDITE could be implemented in platforms such as Rydberg atom arrays [113, 23], superconducting qubits [110], or trapped-ion systems [34]. Such realizations would enable direct observation of the mixed-state criticalities predicted here, bridging theory and experiment. Therefore, the MDITE framework may provide a useful platform for designing quantum algorithms for preparing nontrivial mixed states, potentially enabling efficient simulations of open quantum many-body systems.
Acknowledgements
We thank G.-Y. Zhu, B. Miao, Y. Guo, Z. Bi, C. Wang, and Y. Zou for helpful discussions. Zenan Liu thanks the China Postdoctoral Science Foundation under Grants No.2024M762935 and NSFC Special Fund for Theoretical Physics under Grants No.12447119. Zhe Wang thanks the China Postdoctoral Science Foundation under Grants No.2024M752898. This project is supported by the Scientific Research Project (No.WU2025B011) and the Start-up Funding of Westlake University. The authors thank the high-performance computing center of Westlake University for providing HPC resources.
Data availability. Data are available on Zenodo [21].
Appendix A A brief review of the stochastic series expansion quantum Monte Carlo
The stochastic series expansion (SSE) method is a versatile and efficient quantum Monte Carlo approach for investigating both finite-temperature and ground-state properties of spin and bosonic systems. It reformulates the partition function $Z=\text{Tr}(e^{-\beta H})$ by performing a Taylor expansion of the exponential operator. The expansion is carried out in a chosen basis $\{|\alpha\rangle\}$, enabling efficient stochastic sampling of physical observables.
By writing the Hamiltonian as a sum of local operators,
$H=-\sum_{a,b}H_{a,b},$ (26)
where $a$ refers to the type of operators (diagonal or off-diagonal operator; site or bond operator) and $b$ marks the spatial indices (e.g., the labels of sites or bonds), we have
$Z=$ $\sum_{\alpha}\langle\alpha|e^{-\beta H}|\alpha\rangle$
$=$ $\sum_{\alpha}\sum_{\{H_{a,b}\}}\frac{\beta^{n}(M-n)!}{M!}\langle\alpha|\prod^{M}_{i=1}H_{a(i),b(i)}|\alpha\rangle,$ (27)
where $M$ is the truncation of the expansion order. To ensure that all operator sequences $\prod_{i=1}^{M}H_{a(i),b(i)}$ have the same fixed length $M$, we have introduced a null operator $H_{-1,-1}$ serves as an identity element filling the unoccupied positions in each sequence.
In the SSE method, each Monte Carlo step consists of a diagonal update and an off-diagonal update. The diagonal update alternates between diagonal operators and null operators while preserving the detailed balance condition. Through this process, the total number of non-null operators $n$ in the operator sequence can change.
  1. (a)
    For each null operator, we try inserting a new diagonal operator at a random position with probability
    $P_{\mathrm{add}}(n\rightarrow n+1)=\min\bigg\{\frac{\beta N_{b}\langle\alpha|H_{b}|\alpha\rangle}{M-n},1\bigg\},$ (28)
    where $N_{b}$ denotes the total number of spatial bonds (or sites) available for insertion.
  2. (b)
    For each diagonal operator, we remove it with probability
    $P_{\mathrm{remove}}(n+1\rightarrow n)=\min\bigg\{\frac{M-n+1}{\beta N_{b}\langle\alpha|H_{b}|\alpha\rangle},1\bigg\}.$ (29)
The off-diagonal update will change the number of operators and update between diagonal operators and off-diagonal operators. To improve the update efficiency, some nonlocal update scheme such as the loop update, cluster update, or directed loop update are typically used [81, 80, 92].
Next, we take the TFIM in Sec. IV as an example to illustrate the SSE method, and it is similar to discuss the CDHM in Sec. VI. The Hamiltonian of the TFIM is given by
$H=-J\sum_{\langle ij\rangle}Z_{i}Z_{j}-h\sum_{i}X_{i}.$ (30)
To express this Hamiltonian in the form of Eq. (26), we introduce the following operators:
$H_{0,i}$ $\equiv hI_{i},$
$H_{1,i}$ $\equiv hX_{i},$
$H_{2,\langle ij\rangle}$ $\equiv J(Z_{i}Z_{j}+I_{i}I_{j}).$ (31)
In the computational basis, defined as the eigenbasis of the Pauli-$Z$ operators, $H_{0,i}$ represents a diagonal on-site operator associated with site $i$, $H_{1,i}$ is an off-diagonal on-site operator, and $H_{2,\langle ij\rangle}$ is a diagonal bond operator associated with the site pair $\langle ij\rangle$. The nonzero matrix elements of the operators defined in Eq. (A) are
$\langle\uparrow_{i}|H_{0,i}|\uparrow_{i}\rangle$ $=\langle\downarrow_{i}|H_{0,i}|\downarrow_{i}\rangle=h$
$\langle\uparrow_{i}|H_{1,i}|\downarrow_{i}\rangle$ $=\langle\downarrow_{i}|H_{1,i}|\uparrow_{i}\rangle=h$
$\langle\uparrow_{i}\uparrow_{j}|H_{2,\langle ij\rangle}|\uparrow_{i}\uparrow_{j}\rangle$ $=\langle\downarrow_{i}\downarrow_{j}|H_{2,\langle ij\rangle}|\downarrow_{i}\downarrow_{j}\rangle=2J$ (32)
[…]
Figure 13: The diagram for cluster update in the TFIM model. White rectangle and blue rectangle denotes the diagonal operator and site operator respectively. The yellow area represents a constructed cluster, with the purple solid line indicating the update line. The line stops upon encountering a site operator; when it meets a bond operator, it continues to extend branches from the other legs of that operator. The dashed lines represent periodic boundary conditions in imaginary time.
For the cluster update, we construct clusters according to the following two rules: (i) A cluster is terminated by a site operator; (ii) The update line continues to extend from the other legs of a bond operator, while each bond operator belongs to exactly one cluster. The process is repeated until all clusters are identified. An illustration of the cluster construction is shown in Fig. 13, where the yellow area is a cluster. Finally, each cluster is flipped with probability $1/2$. As we will show below, this update can be readily adapted to the simulation of the extended ensemble in Eq. (9).
Appendix B Quantum Monte Carlo algorithm for simulating the generalized partition function $\mathcal{Q}_{n_{d}}$
We still take the TFIM as an example. Compared with the standard partition function $Z$, the simulation of $\mathcal{Q}_{n_{d}}$ involves multiple replicas and probabilistic boundary conditions that connect them. Fig. 14 shows the SSE configuration of $\bar{\rho}_{1}\propto e^{-\tau H}$, whose QMC simulation is equivalent to that of a Gibbs state with partition function $Z=\text{Tr}(e^{-\tau H})$.
[…]
Figure 14: The measurement-averaged state $\overline{\rho}_{1}$ corresponds to the sampling configuration in the SSE stimulation. In the diagram, white rectangles denote diagonal operators, blue rectangles represent site operators, and red rectangles indicate auxiliary identity operators. Additional auxiliary identity operators (red rectangles) are introduced at each local site. Those auxiliary operators do not affect the physical properties of the system, and they facilitate the treatment of boundary conditions in simulations practically.
Figure 15: The measurement-averaged state $\overline{\rho}_{1}$ corresponds to the sampling configuration in the SSE stimulation, taking two time steps as an example. White rectangle and blue rectangle represents the diagonal operator and site operator respectively, while red rectangle denotes the auxiliary identity operators. In (a), the simulation performs a measurement operation after a time interval of $\tau/2$. At this moment, the manifold closes, and the two time steps are adhered together. In (b), no measurement is performed at $\tau/2$, which the manifold remains open, and the two circuits form a single larger layer. The dashed lines represent the periodic boundary conditions in imaginary time. The two subfigures correspond respectively to the two extreme cases of measurement rates $p=1$ and $p=0$.
Similarly, when the measurement rate is $p=1$, as discussed in Sec. III.2, the SSE configuration related to $\text{Tr}(\rho_{2})$ is illustrated in Fig. 15(a). Since the state undergoes two evolutions, the diagram contains two replicas of $e^{-\tau H}$. To perform the QMC simulations, we note that the boundary conditions must be consistent with the diagrammatic representation introduced in Sec. III, which imposes constraints on the spin configurations near the boundaries of each replica. Since the diagonal update in SSE is local and does not alter the spin states, it can be implemented in the same manner as that in a standard partiton function $Z$. For the nonlocal update, the boundary conditions require that the spins at the interfaces between different replicas remain consistent in Fig. 15(a). Consequently, when a cluster extends across an interface, it may continue into another replica. In practice, this is implemented using auxiliary identity operators: when two replicas are connected, their corresponding auxiliary identity operators are required to belong to the same cluster.
When the measurement rate is zero, the ensemble reduces to the standard partition function with a doubled inverse temperature, as illustrated in Fig. 15(b). In this case, the auxiliary operators are not involved in the cluster update.
As discussed in Sec. III.4, for a general measurement rate $p$, we only need to switch the simulated ensemble between Fig. 14(a) and (b) when the number of time steps is two. Specifically, before each Monte Carlo step, we consider the following merge-split process:
  1. (1)
    If two spins near the boundary in the corresponding replicas are in the same state but are not connected in the diagrammatic representation, they are merged with probability
    $P_{\mathrm{merge}}=\min\bigg\{\frac{p}{1-p},1\bigg\};$ (33)
  2. (2)
    If two spins are in the same state and already connected in the diagrammatic representation, they are split with probability
    $P_{\mathrm{split}}=\min\bigg\{\frac{1-p}{p},1\bigg\}.$ (34)
The merge-split process can be naturally generalized to the case when the number of time steps is greater than two, i.e., more replicas of $e^{-\tau H}$, following the dicussions in Sec. III.5. For the Heisenberg model, the diagonal update, merge–split process, and nonlocal update are implemented in a manner analogous to those used for the TFIM under measurement. The cluster update scheme can be employed to perform updates between diagonal and off-diagonal operators in the Heisenberg model.
Appendix C Continuous-time limit of the 1D model
In this appendix, we perturbatively analyze the limit that $\tau\to 0$, $\tilde{\tau}\sim O(1)$, and $h\sim O(1)$. In this limit, Eq. (18) can be simplified using the BCH formula,
$e^{\tau A}e^{B}=e^{B+\tau A+\frac{\tau}{2}[A,B]+\frac{\tau}{12}[B,[B,A]]+O(\tau^{2})},$ (35)
where we identify
$A\equiv-\tilde{H},\qquad B\equiv\tilde{\tau}\sum_{i}Z_{i}^{L}Z_{i}^{R}.$ (36)
By dropping terms of order $O(\tau^{2})$ and higher, we have
$e^{\tau A}e^{B}\approx e^{-H_{\text{eff}}},$ (37)
with the effective Hamiltonian
$H_{\text{eff}}\equiv-B-\tau A-\frac{\tau}{2}[A,B]-\frac{\tau}{12}[B,[B,A]].$ (38)
For the two-copy TFIM Hamiltonian,
$\tilde{H}=-\frac{1}{2}\sum_{\langle ij\rangle}(Z_{i}^{L}Z_{j}^{L}+Z_{i}^{R}Z_{j}^{R})-\frac{h}{2}\sum_{i}(X_{i}^{L}+X_{i}^{R}),$ (39)
one finds, after straightforward algebra,
$A,B]=-\tilde{\tau}h\sum_{i}\bigg(iY_{i}^{L}Z_{i}^{R}+iZ_{i}^{L}Y_{i}^{R}\bigg),$ (40)
and
$B,[B,A]]=2\tilde{\tau}^{2}h\sum_{i}(X_{i}^{L}+X_{i}^{R}).$ (41)
Substituting these results into Eq. (38), we may rewrite the effective Hamiltonian, up to an overall rescaling by $\tilde{\tau}$, as
$H_{\text{eff}}=H_{Z}+(H_{ZZ}+H_{X}+H_{YZ}),$ (42)
where
$H_{Z}\equiv$ $-\sum_{i}Z_{i}^{L}Z_{i}^{R},$ (43)
$H_{ZZ}\equiv$ $-\alpha_{ZZ}\sum_{\langle ij\rangle}(Z_{i}^{L}Z_{j}^{L}+Z_{i}^{R}Z_{j}^{R}),$ (44)
$H_{X}\equiv$ $-\alpha_{X}\sum_{i}(X_{i}^{L}+X_{i}^{R}),$ (45)
$H_{YZ}\equiv$ $\alpha_{YZ}\sum_{i}(iY_{i}^{L}Z_{i}^{R}+iZ_{i}^{L}Y_{i}^{R}).$ (46)
The couplings are given by
$\alpha_{ZZ}\equiv\frac{\tau}{2\tilde{\tau}},\quad\alpha_{X}\equiv\frac{\tau}{\tilde{\tau}}\bigg(\frac{h}{2}+\frac{\tilde{\tau}^{2}h}{6}\bigg),\quad\alpha_{YZ}\equiv\frac{\tau h}{2}.$ (47)
Moreover, the effective Hamiltonian can be decomposed as
$H_{\mathrm{eff}}=H_{0}+H_{1},$ (48)
where $H_{0}\equiv H_{Z}$ and
$H_{1}\equiv\frac{\tau}{\tilde{\tau}}(H_{ZZ}+H_{X}+H_{YZ})$ (49)
is treated as a perturbation. We remark that, in this case, taking $\tau\to 0$ does not imply turning off the dynamics, but instead corresponds to a Trotterization of the combined evolution and measurement processes.
The unperturbed Hamiltonian $H_{0}$, whose ground-state energy is $E_{0}=-N$, dominates the low-energy Hilbert space and enforces the constraint $Z_{i}^{L}Z_{i}^{R}=+1$ for all $i$. Moreover, a single constraint violation at some site $j$, characterized by $Z_{j}^{L}Z_{j}^{R}=-1$, costs an energy gap $\Delta=2$. In the limit $\tau/\tilde{\tau}\to 0$, transitions out of the low-energy subspace induced by $H_{1}$ are perturbatively suppressed.
Next, we derive the low-energy Hamiltonian of $H_{\text{eff}}$. For convenience, we introduce the projectors
$P\equiv\prod_{i}P_{i},\qquad Q\equiv 1-P,$ (50)
where
$P_{i}\equiv\frac{1+Z_{i}^{L}Z_{i}^{R}}{2}$ (51)
projects onto the subspace with locally aligned left and right spins. Applying a Schrieffer-Wolff (SW) transformation, the low-energy effective Hamiltonian up to second order reads
$H_{\mathrm{eff}}\approx PH_{0}P+H^{(1)}_{\text{eff}}+H^{(2)}_{\text{eff}}$, (52)
where
$H^{(1)}_{\text{eff}}$= $PH_{1}P,$ (53)
$H^{(2)}_{\text{eff}}$= $-PH_{1}Q\frac{1}{H_{0}-E_{0}}QH_{1}P.$ (54)
Since $H_{0}$ imposes a hard constraint, the low-energy physics is governed by the remaining two correction terms.
The first-order term is given by
$H^{(1)}_{\mathrm{eff}}=H_{ZZ}=-\alpha_{ZZ}\sum_{\langle ij\rangle}(Z_{i}^{L}Z_{j}^{L}+Z_{i}^{R}Z_{j}^{R}),$ (55)
which describes FM Ising interactions within each copy.
At the second order, we have
$H^{(2)}_{\mathrm{eff}}=-(\alpha_{X}^{2}-\alpha_{YZ}^{2})\sum_{i}X_{i}^{L}X_{i}^{R},$ (56)
up to an unimportant constant energy shift.
Collecting the leading contributions, the effective low-energy Hamiltonian within the constrained subspace $Z_{i}^{L}Z_{i}^{R}=+1$ takes the form
$H^{\prime}_{\mathrm{eff}}=-\tilde{J}\sum_{\langle ij\rangle}(Z_{i}^{L}Z_{j}^{L}+Z_{i}^{R}Z_{j}^{R})-\tilde{h}\sum_{i}X_{i}^{L}X_{i}^{R},$ (57)
where the effective couplings are given by
$\tilde{J}\equiv\;\alpha_{ZZ}=\frac{\tau}{2\tilde{\tau}}$, (58)
and
$\tilde{h}\equiv\alpha_{X}^{2}-\alpha_{YZ}^{2}=\tau^{2}h^{2}\bigg[\frac{1}{\tilde{\tau}^{2}}\bigg(\frac{1}{2}+\frac{\tilde{\tau}^{2}}{6}\bigg)^{2}-\frac{1}{4}\bigg]>0.$ (59)
Inside the constrained subspace $Z_{i}^{L}Z_{i}^{R}=+1$, the constrained Hilbert space is exactly that of a single qubit per site. Therefore, by defining effective Pauli operators
$\tilde{Z}_{i}\equiv\tilde{Z}_{i}^{L}\equiv\tilde{Z}_{i}^{R},\quad\tilde{X}_{i}\equiv X_{i}^{L}X_{i}^{R},$ (60)
which satisfy the standard Pauli algebra, we have
$H^{\prime}_{\mathrm{eff}}=-2\tilde{J}\sum_{\langle ij\rangle}\tilde{Z}_{i}\tilde{Z}_{j}-\tilde{h}\sum_{i}\tilde{X}_{i}.$ (61)
This Hamiltonian is precisely a single-copy TFIM. For the 1D case, it has a quantum critical point at $2\tilde{J}=\tilde{h}$ according to the Kramers-Wannier duality, which belongs to the 2D classical Ising universality class.
Furthermore,
$2\tilde{J}\gtrless\tilde{h}\quad\Longleftrightarrow\quad\frac{\tau}{\tilde{\tau}}$ $\gtrless\tau^{2}h^{2}\left[\frac{1}{\tilde{\tau}^{2}}\left(\frac{1}{2}+\frac{\tilde{\tau}^{2}}{6}\right)^{2}-\frac{1}{4}\right]$ (62)
[…]
Since $\tau\to 0$, the left-hand side of Eq. (62) becomes parametrically large. Consequently, by increasing $\tilde{\tau}$, or equivalently the measurement rate $p$, the system can be driven across the critical point, inducing a quantum phase transition from the FM ordered phase to the PM disordered phase.
This behavior is physically intuitive. In the continuous time limit, when the measurement rate is much larger than the ITE scale, i.e., $\tilde{\tau}\gg\tau$, the steady state is dominated by the decoherence generated by $e^{\tilde{\tau}\sum_{i}Z_{i}^{L}Z_{i}^{R}}$ in Eq. (18). In this regime, the measurements strongly constrain the system, preventing the ITE from establishing long-range order, even $h<1$.
The same conclusion follows directly from the effective Hamiltonian in Eq. (42). In the extreme limit $\tau=0$, the dynamics is entirely governed by the measurement process, and the steady state coincides with the ground state of $H_{\mathrm{eff}}=-\sum_{i}Z_{i}^{L}Z_{i}^{R}$. This state enforces local alignment between the two copies at each site but does not generate spatial Ising order within either copy.
We emphasize that this mechanism is distinct from that observed in Sec. IV away from the limit, where increasing the measurement rate $p$ instead drives a phase transition from a disordered phase to an ordered phase.
Appendix D Weak-measurement limit of the 1D model
In the weak-measurement limit, $\tilde{\tau}\to 0\sim p/2$, with $\tau\sim O(1)$ and $h\sim O(1)$, one can similarly derive an effective two-copy Hamiltonian
$H_{\text{eff}}=H_{0}+H_{1},$ (63)
where
$H_{0}\equiv\tilde{H}=-\frac{1}{2}\sum_{\langle ij\rangle}(Z_{i}^{L}Z_{j}^{L}+Z_{i}^{R}Z_{j}^{R})-\frac{h}{2}\sum_{i}(X_{i}^{L}+X_{i}^{R}),$ (64)
and
$H_{1}\equiv H_{Z}+H_{YZ}+H_{XZZ}+H_{Y}$ (65)
with
$H_{Z}\equiv$ $\bigg(\frac{p\tau h}{12}-\frac{p}{2\tau}\bigg)\sum_{i}Z_{i}^{L}Z_{i}^{R},$ (66)
$H_{YZ}\equiv$ $-\frac{ph}{4}\sum_{i}\bigg(iY_{i}^{L}Z_{i}^{R}+iZ_{i}^{L}Y_{i}^{R}\bigg),$ (67)
$H_{XZZ}\equiv$ $-\frac{p\tau h}{24}\sum_{\langle ij\rangle}(X_{i}^{L}Z_{j}^{L}Z_{i}^{R}$
$+Z_{i}^{L}X_{j}^{L}Z_{j}^{R}+Z_{i}^{L}X_{i}^{R}Z_{j}^{R}+Z_{j}^{L}Z_{i}^{R}X_{j}^{R},)$ (68)
$H_{Y}\equiv$ $-\frac{p\tau h}{12}\sum_{i}Y_{i}^{L}Y_{i}^{R}.$ (69)
The unperturbed Hamiltonian $H_{0}$ has a global $\mathbb{Z}_{2}^{L}\times\mathbb{Z}_{2}^{R}$ symmetry. Accordingly, the projector onto the ground-state sector can be written as
$P=\prod_{i}P_{i},\quad P_{i}\equiv\frac{1+X_{i}^{L}}{2}\frac{1+X_{i}^{R}}{2}.$ (70)
[…]
This perturbative stability originates not only from the finite excitation gap, but also from the product structure of the ground-state projector $P=\prod_{i}P_{i}$, which imposes simultaneous constraints on the two copies. As a result, any local perturbation necessarily creates gapped excitations and cannot generate nontrivial operators within the low-energy subspace at these orders.
Therefore, the weak-measurement limit also does not exhibit physics that is qualitatively consistent with the numerical results presented in Sec. IV.
Appendix E Cluster formations and long-range orders
In this appendix, we provide an intuitive, QMC-based picture for how local projective measurements enhance collective behavior in the system, ultimately generating a mixed ordered phase.
As introduced in Appendix. A, the nonlocal cluster update is a crucial step to ensure the ergodicity of the QMC algorithm for the TFIM. Therefore, the size of the automatically formed clusters reflects the degree of collective behavior among spins: large clusters emerge when many spins fluctuate coherently. For example, in the 1D TFIM, when the transverse field is very strong $(h\to\infty)$, spins are dominated by the external field, and cluster updates produce only small, local clusters. Conversely, when the field is very weak, $(h\to 0)$, the Ising interactions dominate, connecting all sites into a single large clusters. In this regime, cluster construction naturally produces system-spanning clusters, reflecting the collective alignment of spins in the FM phase.
According to the merge-split process introduced in Appendix. B, the presence of measurements allows update lines to connect across replicas, effectively merging clusters from different replicas. Therefore, starting with the PM ground state of the TFIM, even local measurements can induce the formation of larger clusters during the QMC updates, which is similar to increasing the Ising coupling strength for the ground state of the TFIM. This reflects why measurements can drive the system into a mixed-ordered phase with long-range order, from the perspective of cluster formations in QMC simulations. Similar reasoning also applies to the CDHM, where a cluster update is also used.
We emphasize that for general Hamiltonians, the specific form of the nonlocal update scheme may differ from the cluster update employed in the TFIM and CDHM. Depending on the model, the corresponding update mechanism could also reveal new insights into the intrinsic properties of the states in the MDITE process.
Appendix F Fittings of the other two dynamical exponents for the 1D TFIM
In this appendix we present additional numerical results for the extraction of the dynamical exponent $z$ in the example of 1D TFIM in Sec. IV.5, using Eq. (12). These results complement the analysis presented in Sec. IV.5 and further support the conclusion that the mixed-state critical points exhibit a dynamical exponent close to $z=1$. See Table 4.
$(\tau,h,p)$ $L_{\min}$ $a$ $z$
(1.2,1.8373,0.8) 64 9.2(2) 1.013(5)
72 9.1(3) 1.010(6)
80 9.0(3) 1.009(7)
88 8.6(4) 0.997(9)
96 8.1(5) 0.99(1)
(0.2651,2.5,0.5) 64 2.2(1) 0.97(1)
72 2.2(1) 0.97(1)
80 2.0(1) 0.96(1)
88 1.8(2) 0.93(2)
96 2.0(2) 0.96(3)
Table 4: Extraction of the dynamical exponent $z$ from the relaxation time for two critical points in the example of TFIM. System sizes range from L=64,72,,128L=64,72,...,128. To test the stability of the extracted exponent, small system sizes are gradually discarded. $L_{\min}$ denotes the smallest system size included in the fits.
Appendix G Dynamical scaling of the example of 2D CDHM
Figure 16: When setting $(\tau,g,p)=(1,3.5,0.3538)$ for the MDITE with the 2D CDHM: (a) The relaxation process of $\langle|m|\rangle$ with increasing the number of time steps $n_{d}$ for various system sizes $L$. It satisfies the formula $m=a\exp(-n_{d}/\xi_{\tau})+b$, where $a$, $\xi_{\tau}$ and $b$ are fitting parameters. (b) Extracting the dynamical exponent $z$ from $\xi_{\tau}\sim L^{z}$ with various system size $L$.
Similar to the example of 1D TFIM, we extract the dynamical exponent $z$ for the 2D CDHM by simulating the relaxation dynamics of the order parameter $\langle|m_{z}|\rangle$.
As a representative case, for $(\tau,g)=(1,3.5)$ the phase transition occurs at $p_{c}\approx 0.3538$, as shown in Fig. 12(a) and (b). From the relaxation behavior illustrated in Fig. 16, we extract the relaxation time $\xi_{\tau}$ and obtain a dynamical exponent $z\approx 0.963$, which is close to unity. The stability analysis further confirms $z\approx 1$, as summarized in Table 5.
We also determine the dynamical exponent for two additional parameter sets, $(\tau,g,p)=(2,2.806,0.3)$ and $(0.4679,3.0,0.1)$. The extracted values again yield $z\approx 1$.
$(\tau,g,p)$ $L_{\min}$ $a$ $z$
(1,3.5,0.3538) 12 4.63(6) 0.963(4)
16 4.56(7) 0.959(4)
20 4.64(8) 0.963(5)
24 4.6(1) 0.963(6)
(2,2.806,0.3) 8 11.3(3) 1.092(6)
12 11.3(3) 1.093(6)
16 11.2(3) 1.091(7)
20 11.4(3) 1.096(8)
(0.4679,3.0,0.1) 12 2.17(6) 1.015(8)
16 2.17(6) 1.015(8)
20 2.14(7) 1.011(8)
24 2.08(8) 1.00(1)
Table 5: Extraction of the dynamical exponent $z$ from the relaxation time in the CDHM using system sizes $L\in[8,48]$. Different parameter sets $(\tau,g,p)$ correspond to three quantum critical points. To test the stability of the fits, small system sizes are gradually discarded. The resulting values are consistently close to $z\approx 1$.
Appendix H Results on the absence of off-diagonal long-range order in the 2D CDHM
Figure 17: Finite size extrapolation of off-diagonal spin-spin correlation $C_{XY}(L/2,L/2)$ at different measurement probability $p$ with (a) $\tau=16$ and (b) $\tau=100$ .
In this appendix, we present additional numerical results demonstrating the absence of off-diagonal long-range order in the 2D CDHM under the MDITE protocol, as discussed in Sec. VI. Combining the algorithm in Sec. E and the QMC tomography method [57, 99, 12], we compute the correlation function
$C_{XY}(\textbf{r})=\langle X_{\textbf{r}}X_{0}+Y_{\textbf{r}}Y_{0}\rangle.$ (71)
To extrapolate to the limit $|\textbf{r}|\to\infty$, we evaluate $C_{XY}(\textbf{r})$ at the maximum separation $\textbf{r}=(L/2,L/2)$ for each system size $L\times L$. We perform a finite-size extrapolation of $C_{XY}(L/2,L/2)$ as a function of $1/L$. As shown in Fig. 17, for sufficiently long imaginary time $\tau$, $C_{XY}(L/2,L/2)$ extrapolates to zero in the thermodynamic limit ($L\to\infty$), indicating the absence of off-diagonal long-range order in the mixed states generated by the MDITE protocol.